### v1

# Contrastive Disentangling: Fine-grained representation learning through multi-level contrastive learning without class priors

Abstract
Recent advancements in unsupervised representation learning often leverage class information to enhance feature extraction and clustering performance. However, this reliance on class priors limits the applicability of such methods in real-world scenarios where class information is unavailable or ambiguous. In this paper, we propose Contrastive Disentangling (CD), a simple and effective framework that learns representations without any reliance on class priors. Our framework employs a multi-level contrastive learning strategy that combines instance-level and feature-level losses with a normalized entropy loss to learn semantically rich and fine-grained representations. Specifically, (1) the instance-level contrastive loss encourages the separation of feature representations for different samples, (2) the feature-level contrastive loss promotes independence among the feature head predictions, and (3) the normalized entropy loss encourages the feature heads to capture meaningful and prevalent attributes from the data. These components work together to enable CD to significantly outperform existing methods, as demonstrated by extensive experiments on benchmark datasets including CIFAR-10, CIFAR-100, STL-10, and ImageNet-10, particularly in scenarios where class priors are absent. The code is available at https://github.com/Hoper-J/Contrastive-Disentangling.
keywords:
Disentangling representations , Clustering , Contrastive learning , Representation learning , Unsupervised learning
[…]
[fafu]organization=School of computer science and information engineering, addressline=Fujian Agriculture and Forestry University, city=Fuzhou, postcode=350002, country=China
1 Introduction
The growing importance of unsupervised learning in processing large-scale unlabeled datasets has increased the demand for effective feature extraction techniques that do not rely on predefined class labels. In computer vision tasks, unsupervised representation learning has become a powerful solution for deriving meaningful feature representations from raw data. A key advancement in this domain is contrastive learning, which has emerged as a prominent paradigm within unsupervised representation learning. By maximizing consistency between different augmented views of the same data instance while minimizing similarity between different instances, contrastive learning has significantly advanced feature extraction (Ye et al., 2019; Chen et al., 2020).
As research in this field progresses, numerous contrastive learning methods (Van Gansbeke et al., 2020; Li et al., 2021; Qian, 2023) have started to utilize class number information to further improve performance. These approaches use class numbers as prior knowledge during network construction, achieving state-of-the-art results in unsupervised clustering tasks. This suggests that incorporating class information during training yield more discriminative feature representations. However, in practical applications, class information is often unavailable, or the data may lack clear structure, limiting the applicability of these methods.
To address this limitation, we propose a novel framework, Contrastive Disentangling (CD), which enhances feature extraction and disentanglement in unsupervised learning tasks through a multi-level contrastive learning strategy and an innovative disentanglement loss function, all without relying on class number information. Unlike methods that rely on additional complex modules, such as self-labeling, our approach focuses on validating the effectiveness of the core framework to ensure its broad applicability.
Figure 1: Multi-level Contrastive Learning in the CD Framework. Instance-level (top) operates on image instances, while feature-level (bottom) focuses on features visualized with LIME.
As illustrated in Figure 1, the CD framework incorporates two levels of contrastive learning: instance-level and feature-level. Instance-level contrastive learning aims to distinguish different image instances within the feature space, while feature-level contrastive learning seeks to differentiate the features captured by distinct prediction heads. In the feature-level module, we employ the LIME (Local Interpretable Model-Agnostic Explanations) technique (Ribeiro et al., 2016) to interpret the outputs of specific feature prediction heads, providing an intuitive visualization of the diverse feature attributes captured by the model across the image batch. In summary, the main contributions of this paper include:
  1. 1.
    We introduce feature prediction heads, a novel addition to the feature learning process. Using normalized entropy loss, these heads effectively capture common features prevalent in the dataset, thereby enhancing the representational capacity and generalization of the features.
  2. 2.
    A feature-level contrastive learning strategy is proposed, ensuring that different feature prediction heads focus on learning distinct and complementary features, thus promoting feature disentanglement.
  3. 3.
    We propose a simple and easy-to-use framework that significantly enhances the learning of fine-grained features in unsupervised learning tasks without relying on class information. Extensive experiments across multiple benchmark datasets demonstrate the superior performance of the CD framework, particularly in metrics like NMI, where it outperforms compared methods. Through comprehensive ablation studies, we validate the effectiveness of each component, providing new insights in unsupervised representation learning.
2 Related Work
2.1 Unsupervised Representation Learning
Unsupervised representation learning seeks to extract meaningful features from unlabeled data. The field is primarily dominated by two approaches: generative models and contrastive learning. Generative models (Bengio et al., 2006; Kingma, 2013; Goodfellow et al., 2014) learn to generate samples resembling the true data distribution by modeling the underlying latent space. However, these models often struggle to learn invariant feature representations, which are insensitive to certain input transformations like rotation, translation, and scaling (Ntelemis et al., 2023).
To overcome these challenges, instance-level contrastive learning methods (He et al., 2020; Chen et al., 2020; Grill et al., 2020; Chen and He, 2021) have been proposed. These approaches significantly enhance the model’s capacity to extract invariant features by maximizing the similarity between different views of the same instance. These methods have exhibited superior performance across various downstream tasks, underscoring their efficacy in unsupervised learning scenarios.
2.2 Utilization of Class Number Information
Recent approaches have leveraged class number information as additional prior knowledge to enhance model performance. Methods (Van Gansbeke et al., 2020; Li et al., 2021; Niu et al., 2022; Qian, 2023) incorporate class number to constrain the final output during network construction. Several of these methods (Van Gansbeke et al., 2020; Niu et al., 2022; Qian, 2023) also employ pseudo-labeling techniques, further enhancing the model’s performance in unsupervised clustering tasks. These advancements indicate that class-level information can effectively assist models in distilling more semantically meaningful features.
However, this dependence on class information constrains the applicability of these methods in real-world settings where class labels or the number of classes are frequently unknown or ambiguous. This limitation emphasizes the necessity for more adaptive approaches that can perform effectively without relying on class priors.
2.3 Multi-level Contrastive Learning
The introduction of cluster-level contrastive learning by CC (Li et al., 2021) for clustering tasks has unveiled new avenues in contrastive learning research. However, partitioning based on predefined categories may not adequately capture the fine-grained features inherent in datasets, potentially constraining the model’s feature extraction capabilities. Building upon this observation, we propose a novel architecture that leverages feature space representations as inputs to feature prediction heads. By extracting common features and implementing feature-level contrastive learning, our approach achieves relative decoupling of distinct feature heads. This facilitates the capture of more discriminative and fine-grained features, thereby advancing the state-of-the-art in unsupervised representation learning.
3 Method
3.1 Proposed Contrastive Disentangling Framework
The proposed CD framework leverages multi-level contrastive learning, as illustrated in Figure 2. It consists of the following key components:
Figure 2: Architecture of the proposed Contrastive Disentangling (CD) framework. The top section illustrates the shared backbone network, instance-level projector, and feature-level predictor, while the bottom section displays the optimization objectives for instance-level and feature-level contrastive learning.
  • 1.
    Data Augmentation. Each input batch $x$ undergoes a series of data augmentation operations, such as random cropping, horizontal flipping, color distortion, and blurring. These augmentations generate two correlated views, denoted as $x^{(1)}$ and $x^{(2)}$, enabling the model to focus on semantically relevant features that remain invariant under these transformations.
  • 2.
    Feature Extraction via ResNet Backbone. A ResNet architecture (He et al., 2016) is used as the shared backbone network $f(\cdot)$, which extracts feature representations from the augmented views $x^{(1)}$ and $x^{(2)}$. This choice aligns with prevailing practices in contrastive learning (Chen et al., 2020; Li et al., 2021). We denote the extracted feature representations as $h^{(1)}=f(x^{(1)})$ and $h^{(2)}=f(x^{(2)})$.
  • 3.
    Instance Projector. The feature representations $h^{(1)}$ and $h^{(2)}$ are passed through an instance projector $g(\cdot)$, implemented as a multi-layer perceptron (MLP). This MLP consists of two linear layers, batch normalization (Ioffe, 2015), and a ReLU activation function. The output is a set of latent representations $z^{(1)}=g(h^{(1)})$ and $z^{(2)}=g(h^{(2)})$, which are crucial for the contrastive learning process (Chen et al., 2020).
  • 4.
    Feature Predictor. The latent representations $z^{(1)}$ and $z^{(2)}$ are fed into a feature predictor $h(\cdot)$, structured similarly to the instance projector but using sigmoid activation in the final layer. The feature predictor outputs predictions $y^{(1)}=h(z^{(1)})$ and $y^{(2)}=h(z^{(2)})$, organized into multiple prediction heads. Each head predicts a distinct feature attribute. We will validate the design of the feature predictor in Section 4.5.3.
  • 5.
    Loss Functions. The total loss comprises three components: the instance-level contrastive loss, the feature-level contrastive loss, and an additional normalized entropy loss. These losses work together to enhance feature diversity, promote independence among learned features, and ensure a wide range of feature predictions.
3.2 Instance-level Contrastive Learning
The objective of instance-level contrastive learning is to pull the representations of different augmented views of the same image closer in feature space while pushing apart representations of different images. Specifically, given an input batch $x$, two views $x^{(1)}$ and $x^{(2)}$ are generated via data augmentation, which are then processed by the shared backbone network and instance projector to yield latent representations $z^{(1)}=g(f(x^{(1)}))$ and $z^{(2)}=g(f(x^{(2)}))$.
These latent representations, $z^{(1)}$ and $z^{(2)}$, are concatenated into a matrix $\mathbf{z}\in\mathbb{R}^{2N\times d}$, where $N$ is the batch size and $d$ is the feature dimension:
$\mathbf{z}=\begin{bmatrix}z^{(1)}\\ z^{(2)}\end{bmatrix}=\begin{bmatrix}z_{1}\\ z_{2}\\ \vdots\\ z_{2N}\end{bmatrix}$ (1)
To quantify the similarity between any two latent representations, we introduce the cosine similarity measure $\text{sim}_{z}(i,j)$:
$\text{sim}_{z}(i,j)=\frac{z_{i}\cdot z_{j}}{\|z_{i}\|\|z_{j}\|}$ (2)
where $i,j\in\{1,2,\dots,2N\}$ and $N$ is the batch size. Based on this similarity measure, we optimize the model by maximizing the similarity between different views of the same instance while minimizing the similarity between different instances. The loss function for instance $i$ is defined as:
$\ell_{i}^{\text{inst}}=-\log\frac{\exp(\text{sim}_{z}(i,\text{pos}(i))/\tau_{\text{inst}})}{\sum_{k=1}^{2N}\mathbbm{1}_{[k\neq i]}\exp(\text{sim}_{z}(i,k)/\tau_{\text{inst}})}$ (3)
Here, $\text{pos}(i)$ represents the index of the positive sample corresponding to $i$. If $i\leq N$, then $\text{pos}(i)=i+N$; otherwise, $\text{pos}(i)=i-N$. The temperature parameter $\tau_{\text{inst}}$ within the instance-level contrastive loss governs the smoothness of the similarity distribution. Additionally, the indicator function $\mathbbm{1}_{[k\neq i]}$ assigns a value of 1 when $k\neq i$, thereby excluding the self-similarity of $z_{i}$. The instance-level contrastive loss is then calculated across all instance views:
$\mathcal{L}_{\text{inst}}=\frac{1}{2N}\sum_{i=1}^{2N}\ell_{i}^{\text{inst}}$ (4)
3.3 Disentanglement Loss
To enhance the independence and diversity of the learned features, we introduce two complementary loss functions: feature-level contrastive loss and normalized entropy loss. These losses work together to promote the learning of disentangled and fine-grained feature attributes.
Feature-level Contrastive Loss. The feature predictor $h(\cdot)$ is designed with multiple independent prediction heads, each focused on predicting a fine-grained feature attribute. After processing the input views $x^{(1)}$ and $x^{(2)}$ through the backbone network and instance projector to generate latent representations $z^{(1)}$ and $z^{(2)}$, these representations are fed into the feature predictor $h(\cdot)$, yielding the corresponding feature predictions $y^{(1)}$ and $y^{(2)}$.
To align with the instance-level contrastive loss, these predictions $y$ are transposed and organized into a matrix $\mathbf{y}\in\mathbb{R}^{2K\times N}$, where $K$ denotes the number of features and $N$ represents the batch size:
$\mathbf{y}=\begin{bmatrix}y^{(1)\top}\\ y^{(2)\top}\end{bmatrix}=\begin{bmatrix}y_{1}\\ y_{2}\\ \vdots\\ y_{2K}\end{bmatrix}$ (5)
In this matrix, $y_{i}$ and $y_{\text{pos}(i)}$ correspond to the prediction results of feature head $i$ on the same batch of samples under two different views. We then define the similarity function $\text{sim}_{y}(i,j)$ to measure the cosine similarity between any two feature prediction vectors:
$\text{sim}_{y}(i,j)=\frac{y_{i}\cdot y_{j}}{\|y_{i}\|\|y_{j}\|}$ (6)
To ensure that feature head $i$ learns features different from other heads, the feature-level contrastive loss is defined as follows:
$\ell_{i}^{\text{feat}}=-\log\frac{\exp(\text{sim}_{y}(i,\text{pos}(i))/\tau_{\text{feat}})}{\sum_{k=1}^{2K}\mathbbm{1}_{[k\neq i]}\exp(\text{sim}_{y}(i,k)/\tau_{\text{feat}})}$ (7)
Here, $\tau_{\text{feat}}$ is the temperature parameter for the feature-level contrastive loss, which controls the degree of similarity enforced between different features. The primary objective is to encourage each prediction head to learn disentangled feature attributes. The final feature-level contrastive loss is calculated as follows:
$\mathcal{L}_{\text{feat}}=\frac{1}{2K}\sum_{i=1}^{2K}\ell_{i}^{\text{feat}}$ (8)
Normalized Entropy Loss. To further promote the learning of common features, we introduce a normalized entropy loss $\mathcal{L}_{\text{entropy}}$. This loss encourages the model to produce balanced and varied predictions by maximizing the entropy of the predicted outputs. By doing so, it prevents the model from overfitting to specific features and promotes the capture of more generalizable attributes. The normalized binary entropy $H(y_{i})$ for each feature prediction vector $y_{i}$ is defined as:
$H(y_{i})=-\frac{1}{N\log(2)}\sum_{j=1}^{N}\left(y_{ij}\log(y_{ij})+(1-y_{ij})\log(1-y_{ij})\right)$ (9)
Subsequently, the normalized entropy loss $\mathcal{L}_{\text{entropy}}$ is computed as the mean entropy across all feature predictions:
$\mathcal{L}_{\text{entropy}}=\frac{1}{2K}\sum_{i=1}^{2K}H(y_{i})$ (10)
Finally, the disentanglement loss $\mathcal{L}_{\text{disentanglement}}$ combines the feature-level contrastive loss and the normalized entropy loss:
$\mathcal{L}_{\text{disentanglement}}=\mathcal{L}_{\text{feat}}-\alpha\cdot\mathcal{L}_{\text{entropy}}$ (11)
where $\alpha$ is a hyperparameter that controls the weight of the normalized entropy loss. It balances the learning of common features within the dataset and the disentanglement of independent feature representations. Minimizing this combined loss enables our CD framework to effectively facilitate the unsupervised extraction and disentanglement of meaningful features from the input data. The efficacy of $\mathcal{L}_{\text{disentanglement}}$ is further validated in the experimental section 4.5.3.
Input: Dataset $X$, Batch size $N$, Feature number $K$, Temperature constants $\tau_{\text{inst}}$ and $\tau_{\text{feat}}$
Output: Total loss $\mathcal{L}_{\text{total}}$
for each minibatch $\mathbf{x}$ of size $N$ in $X$ do
      // Apply data augmentation
       $\mathbf{x}\rightarrow\{\mathbf{x}^{(1)},\mathbf{x}^{(2)}\}$;
      for each view $\mathbf{x}^{(v)},v\in\{1,2\}$ do
             // Extract features
             $\mathbf{h}^{(v)}=f(\mathbf{x}^{(v)})$;
            // Project to latent space
             $\mathbf{z}^{(v)}=g(\mathbf{h}^{(v)})$;
            // Generate feature head predictions
             $\mathbf{y}^{(v)}=h(\mathbf{z}^{(v)})$;
       end for
      Combine $z^{(1)}$, $z^{(2)}$ by Eq.(1) and $y^{(1)}$, $y^{(2)}$ by Eq.(5);
      Compute cosine similarities $\text{sim}_{z}(i,j)$, $\text{sim}_{y}(i,j)$ by Eq. (2), (6);
      Calculate instance-level contrastive loss $\mathcal{L}_{\text{inst}}$ by Eq. (3), (4);
      Calculate feature-level contrastive loss $\mathcal{L}_{\text{feat}}$ by Eq. (7), (8);
      Calculate normalized entropy loss $\mathcal{L}_{\text{entropy}}$ by Eq. (9), (10);
      Combine all losses $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{inst}}+\mathcal{L}_{\text{feat}}-\alpha\cdot\mathcal{L}_{\text{entropy}}$;
      Update parameters of $f$, $g$, and $h$ to minimize $\mathcal{L}_{\text{total}}$;
end for
return $\mathcal{L}_{\text{total}}$
Algorithm 1 Contrastive Disentangling Framework
4 Experiment
4.1 Datasets
We conducted our experiments on four widely-used computer vision datasets: CIFAR-10, CIFAR-100 (Krizhevsky and Hinton, 2009), STL-10 (Coates et al., 2011), and ImageNet-10 (Deng et al., 2009). These datasets serve as benchmarks to rigorously evaluate the effectiveness of our proposed Contrastive Disentangling (CD) framework. Table 1 provides an overview of the dataset splits, sample sizes, and the number of classes used.
Table 1: Overview of datasets used for training and evaluation
Dataset Split Samples Classes
CIFAR-10 train+test 60,000 10
CIFAR-100-20 train+test 60,000 20
STL-10 train+test 13,000 10
ImageNet-10 train 13,000 10
For the CIFAR-100 dataset, we used 20 super-classes as ground-truth labels instead of the 100 fine-grained classes, aligning with benchmark settings in previous studies. Furthermore, to ensure the rigor of our architecture evaluation, we did not use any unseen data during training, such as the unlabeled data in STL-10.
4.2 Experimental Settings
Data Augmentation. To ensure fair comparison with existing methods (Li et al., 2021), we adopted the same data augmentation strategies and used ResNet-34 as the backbone network for feature extraction. Specifically, all datasets underwent random cropping to 224x224, random horizontal flipping, color jittering, and grayscale transformation. Notably, Gaussian blur was applied only to the ImageNet-10 dataset, as upscaling images to 224x224 for smaller datasets (e.g., CIFAR-10, CIFAR-100, and STL-10) inherently introduces a similar blurring effect.
Training Details. In our experiments, the hidden dimension and the number of features were both set equal to the batch size. We employed a cosine decay schedule without restarts (Chen et al., 2020; Loshchilov and Hutter, 2016) and gradient clipping to enhance the stability of model training. The effectiveness of these optimizations is demonstrated in the ablation study presented in Table 2. Additionally, no other techniques (e.g., self-labeling) were used to boost model performance. The hyperparameter $\alpha$, which balances the diversity of feature representations, was simply set to 1. The remaining training settings were consistent with prior contrastive learning work (Li et al., 2021): a random seed of 42 was used, and training was conducted for 1000 epochs. The Adam optimizer (Kingma, 2014) was employed, with a learning rate of $3\times 10^{-4}$ and no weight decay. The temperature parameters $\tau_{\text{inst}}$ and $\tau_{\text{feat}}$ were set to 0.5 and 1, respectively.
Environment. The experiments were conducted on an NVIDIA RTX 3090 GPU (24GB) and an 18 vCPU AMD EPYC 9754 128-Core Processor. We experimented with batch sizes of 128 and 256. The training time for the CIFAR-10 and CIFAR-100 datasets was approximately 48 GPU hours, while for the STL-10 and ImageNet-10 datasets, the training time was around 11 GPU hours. To ensure reproducibility and transparency, detailed logs of all experimental procedures are available at here.
Evaluation Metrics. We employed three clustering metrics — Normalized Mutual Information (NMI), Adjusted Rand Index (ARI), and Accuracy (ACC) — to evaluate the feature extraction and disentanglement capabilities of the model.
4.3 Compared Methods
To rigorously assess the performance of our proposed method in various unsupervised learning scenarios, we compared it against a diverse set of representative methods, chosen for their relevance to the challenges addressed by our approach. Alongside the traditional K-means (MacQueen et al., 1967) clustering algorithm, we evaluated several fully unsupervised models—AE (Bengio et al., 2006), DeCNN (Zeiler et al., 2010), DAE (Vincent et al., 2010), VAE (Kingma, 2013), and DCGAN (Radford, 2015)—that span a spectrum of approaches from classical techniques to modern deep generative models. These methods, like our proposed Contrastive Disentangling (CD) model, do not rely on any prior class information during training, including the number of classes.
We also compared our method with unsupervised models that incorporate class number information—such as JULE (Yang et al., 2016), DEC (Xie et al., 2016), DAC (Chang et al., 2017), DCCM (Wu et al., 2019), PICA (Huang et al., 2020), and CC (Li et al., 2021)—to illustrate how our approach performs relative to methods that leverage this additional knowledge. To ensure a fair and consistent comparison, all experiments were conducted using the full set of samples, adhering to identical training and evaluation protocols.
4.4 Experimental Results
We evaluated our model at two key stages: feature extraction and feature disentanglement. Specifically, we applied K-means clustering to both the backbone network output and the final model output to assess the model’s capabilities in feature extraction and disentanglement, respectively.
Feature Extraction. Table 2 presents a comparative analysis of various models at the feature extraction stage, utilizing the NMI, ARI, and ACC metrics. Across four datasets—CIFAR-10, CIFAR-100-20, STL-10, and ImageNet-10—our Contrastive Disentanglement (CD) method, implemented with batch sizes of 128 (CD-128) and 256 (CD-256), demonstrated superior performance compared to other fully unsupervised learning models, including AE, DeCNN, DAE, VAE, and DCGAN. In comparison to the contrastive learning method CC, which utilizes prior information about the number of classes, CD consistently produced the highest NMI scores across all datasets, with notable improvements in STL-10 and ImageNet-10, while maintaining superior overall performance in most other cases.
Table 2: Feature extraction performance comparison across models. The 1st/2nd best results are indicated in bold/underlined
Dataset Metrics K-means AE DeCNN DAE VAE DCGAN CC CD-128 CD-256
[…]
Feature Disentanglement. Table 3 presents a comparison of our method with other unsupervised methods that utilize number of classes information at the final output stage. The results demonstrate that the CD method exhibited excellent performance across all metrics, particularly on the STL-10 and ImageNet-10 datasets. For CIFAR-10 and CIFAR-100-20, although the CD method’s ARI and ACC metrics were slightly lower than those of the CC method, this may be attributed to the CD method not utilizing the number of classes information to constrain the final feature prediction. Notably, on the STL-10 dataset, while the PICA method achieved a slightly higher ACC (0.713) due to the use of additional unlabeled data, the CD-128 model reached a comparable ACC of 0.702 without using extra data, and the CD-256 model further improved this to 0.734.
Table 3: Final output performance comparison across models. The 1st/2nd best results are indicated in bold/underlined, and models that utilized unlabeled datasets during training for STL-10 are marked with
Dataset Metrics JULE DEC DAC DCCM PICA CC CD-128 CD-256
[…]
These results provide strong evidence that the CD method can achieve excellent performance without requiring the number of classes information. It is particularly noteworthy that the CD method exhibited outstanding performance even with smaller batch sizes. For a detailed visualization of the CD-256 model’s feature representations using t-SNE, please see A.
4.5 Ablation Studies
To demonstrate the effectiveness of each component in our model, we conducted a series of ablation studies using the STL-10 and ImageNet-10 datasets with a batch size of 128. In sections 4.5.1 and 4.5.2, we incrementally added modules to validate their effectiveness based on the framework and objective function shown in Figure 2. In sections 4.5.3 and 4.5.4, we removed certain modules to verify their necessity.
4.5.1 Normalized Entropy Loss and Batch Normalization
The impact of normalized entropy loss (NE) and batch normalization (BN) on model performance was initially evaluated, as shown in Table 4. The experimental results demonstrate that both normalized entropy loss and batch normalization significantly enhance model performance, with their synergistic application yielding even more substantial performance improvements. Notably, NE exhibited particularly significant improvements during the feature disentanglement phase, underscoring its pivotal role in facilitating effective feature disentanglement.
Table 4: Effect of normalized entropy loss (NE) and batch normalization (BN) on performance across backbone and final output.
Dataset NE BN Backbone Feature
NMI ARI ACC NMI ARI ACC
STL-10 - - $0.519\pm 0.029$ $0.265\pm 0.037$ $0.538\pm 0.050$ $0.352\pm 0.040$ $0.065\pm 0.028$ $0.354\pm 0.041$
- $0.638\pm 0.021$ $0.482\pm 0.035$ $0.695\pm 0.031$ $0.656\pm 0.005$ $0.527\pm 0.004$ $0.671\pm 0.007$
- $0.659\pm 0.017$ $0.495\pm 0.036$ $0.653\pm 0.037$ $0.367\pm 0.052$ $0.069\pm 0.034$ $0.306\pm 0.044$
$\mathbf{0.677\pm 0.020}$ $\mathbf{0.547\pm 0.031}$ $\mathbf{0.705\pm 0.039}$ $\mathbf{0.669\pm 0.006}$ $\mathbf{0.543\pm 0.016}$ $\mathbf{0.681\pm 0.020}$
ImageNet-10 - - $0.632\pm 0.022$ $0.287\pm 0.041$ $0.620\pm 0.042$ $0.549\pm 0.038$ $0.163\pm 0.037$ $0.424\pm 0.040$
- $0.748\pm 0.020$ $0.532\pm 0.054$ $0.779\pm 0.047$ $0.887\pm 0.002$ $0.861\pm 0.002$ $0.927\pm 0.001$
- $0.837\pm 0.018$ $0.733\pm 0.050$ $0.831\pm 0.055$ $0.538\pm 0.043$ $0.152\pm 0.045$ $0.416\pm 0.055$
$\mathbf{0.857\pm 0.021}$ $\mathbf{0.772\pm 0.068}$ $\mathbf{0.858\pm 0.071}$ $\mathbf{0.890\pm 0.004}$ $\mathbf{0.864\pm 0.006}$ $\mathbf{0.929\pm 0.004}$
To further illustrate the impact of normalized entropy loss on feature disentanglement, a visual representation of the model outputs is presented in Figure 3.
(a) Default
(b) noNEloss
Figure 3: Visualization of transposed model outputs with/without normalized entropy loss (NE loss). Each row corresponds to the prediction of a specific feature attribute across the entire batch. The left visualization (a) shows the output with NE loss enabled, focusing on common features, while the right visualization (b) shows the output without NE loss, capturing more unique, image-specific features.
This visualization clearly demonstrates the effectiveness of NE loss in guiding the model to capture common features across the dataset, rather than focusing on unique features specific to individual images. The rows in Figure 3 represent the feature attributes learned by our model, as depicted in the feature-level part of Figure 1 and further visualized in B. This visual evidence corroborates the quantitative results and emphasizes the significance of NE loss in the proposed methodology.
4.5.2 Training Tricks
Following the validation of core modules, we further evaluated the contribution of cosine decay schedule and gradient clipping training techniques, as shown in Table 5.
Table 5: Effect of cosine decay schedule (Sch.) and gradient clipping (Clip.) on model performance.
Dataset Sch. & Clip. NMI ARI ACC
STL-10 - $0.669\pm 0.006$ $0.543\pm 0.016$ $0.681\pm 0.020$
$\mathbf{0.679\pm 0.007}$ $\mathbf{0.564\pm 0.023}$ $\mathbf{0.720\pm 0.026}$
ImageNet-10 - $0.890\pm 0.004$ $0.864\pm 0.006$ $0.929\pm 0.004$
$\mathbf{0.898\pm 0.003}$ $\mathbf{0.868\pm 0.006}$ $\mathbf{0.931\pm 0.003}$
The results show that both the cosine decay schedule and gradient clipping further enhance the model’s final performance and improve the stability of the results.
4.5.3 Feature Prediction Head
Next, we evaluated the effect of removing the feature prediction head, which directly reduces the model to an instance-level contrastive learning architecture, similar to SimCLR. This experiment was designed to verify the critical role of the disentanglement loss $\mathcal{L}_{\text{disentanglement}}$in the feature disentanglement process.
Table 6: Effect of feature-level head on model performance.
Dataset Feature-Level Head NMI ARI ACC
STL-10 - $0.599\pm 0.029$ $0.458\pm 0.043$ $0.649\pm 0.053$
$\mathbf{0.679\pm 0.007}$ $\mathbf{0.564\pm 0.023}$ $\mathbf{0.720\pm 0.026}$
ImageNet-10 - $0.708\pm 0.031$ $0.413\pm 0.060$ $0.608\pm 0.065$
$\mathbf{0.898\pm 0.003}$ $\mathbf{0.868\pm 0.006}$ $\mathbf{0.931\pm 0.003}$
As shown in Table 6, the results demonstrate that the feature prediction head significantly improves the model’s performance in the final layers (see Table 9 for backbone results). Removing this head leads to a noticeable degradation in performance. This confirms the critical contribution of the feature-level contrastive learning strategy, which enhances the model’s ability to disentangle features across different prediction heads.
4.5.4 Data Augmentation
This section verifies the necessity of the dual-view data augmentation strategy for feature disentanglement.
Table 7: Effect of single-view vs. dual-view data augmentation on model performance.
Dataset $x^{(1)}$ $x^{(2)}$ NMI ARI ACC
STL-10 - $0.643\pm 0.004$ $0.555\pm 0.012$ $0.717\pm 0.014$
$\mathbf{0.679\pm 0.008}$ $\mathbf{0.564\pm 0.023}$ $\mathbf{0.720\pm 0.026}$
ImageNet-10 - $0.874\pm 0.001$ $0.847\pm 0.001$ $0.918\pm 0.001$
$\mathbf{0.898\pm 0.003}$ $\mathbf{0.868\pm 0.006}$ $\mathbf{0.931\pm 0.003}$
As shown in Table 7, the results demonstrate that dual-view augmentation is crucial in the feature disentanglement stage for capturing common features in the dataset. However, in the feature extraction stage, single-view augmentation can achieve satisfactory performance (see Table 10 for details).
5 Conclusion
In this work, we proposed the Contrastive Disentangling (CD) framework, a novel approach for enhancing feature extraction and disentanglement in unsupervised learning without relying on prior knowledge of class numbers. The CD framework integrates multi-level contrastive learning with an innovative disentanglement loss function, emphasizing simplicity and core effectiveness. Experimental results demonstrate that CD consistently outperforms class-dependent methods with simple architectures (i.e., without additional structures such as self-labeling), particularly in NMI, across multiple benchmark datasets including CIFAR-10, CIFAR-100, STL-10, and ImageNet-10. Moreover, CD exhibits robust performance in small-batch settings, effectively capturing common features across diverse data distributions. Comprehensive ablation studies corroborate the efficacy of each framework component, validating its overall effectiveness and potential for wide-scale application in unsupervised learning scenarios.
6 Acknowledgments
This paper was supported by the Fujian Provincial Natural Science Foundation of China (No.2021J01129), the Fujian Provincial Higher Education Technology Research Association Fund Project (No.H2000134A), and the Fujian Agriculture and Forestry University Horizontal Technology Innovation Fund (No.KHF190015).
Appendix A t-SNE Visualization of Backbone and Feature Predictor Outputs
In this appendix, we present the t-SNE (t-Distributed Stochastic Neighbor Embedding) visualizations (Van der Maaten and Hinton, 2008) of the outputs both the backbone feature extraction layer and the feature predictor of our model, across all datasets: CIFAR-10, CIFAR-100, STL-10, and ImageNet-10. These visualizations are shown in Figures 4-7, where (a) represents the backbone output and (b) represents the feature predictor output for each dataset.
(a) Backbone
(b) Feature
Figure 4: t-SNE visualization for CIFAR-10
(a) Backbone
(b) Feature
Figure 5: t-SNE visualization for CIFAR-100
(a) Backbone
(b) Feature
Figure 6: t-SNE visualization for STL-10
(a) Backbone
(b) Feature
Figure 7: t-SNE visualization for ImageNet-10
Appendix B Visualizing Feature Prediction Heads with LIME
We utilize the LIME (Local Interpretable Model-Agnostic Explanations) technique (Ribeiro et al., 2016) to visualize the outputs of the feature prediction heads. LIME is configured with $num\_samples$=5000, meaning that for each image, LIME generates 5000 perturbed samples to create an interpretable explanation. We only visualize images where the feature prediction heads produce outputs greater than 0.75.
Figure 8: Visualization of the attention from five feature prediction heads, with yellow-circled areas highlighting the regions that the heads focus on.
The feature prediction heads visualized in Figure 8 correspond to the indices 1, 5, 6, 17, and 18. These specific heads were selected to ensure a visually meaningful comparison, as some other heads produced fewer than six images that met the 0.75 threshold. The yellow highlighted areas indicate the regions that these heads attend to.
Figure 9 presents the same heads applied to a new batch of images, with a mask applied to remove areas that were not attended to.
Figure 9: Visualization of the same five feature prediction heads on a new batch of images, with non-attended regions masked out.
The model used for this visualization was pre-trained on the ImageNet-10 dataset with the following hyperparameters: batch size=256, epochs=1000, and learning rate=0.0003. All other settings were kept as defaults from our repository, where the pre-trained model files and complete experimental processes for all other experiments in this study are also available.
Appendix C Ablation Studies on Backbone Feature Extraction
This section provides additional ablation studies focused on evaluating the backbone feature extraction capabilities of our proposed framework. These experiments assess the contribution of various framework components by examining their impact on the backbone’s ability to extract meaningful features. Specifically, we analyze the performance of the backbone under different configurations by removing or modifying key components such as the instance projector, feature predictor, and data augmentation strategies.
The experiments are conducted on the STL-10 and ImageNet-10 datasets with the same batch size of 128, as used in the main experiments. We report the results using metrics such as Normalized Mutual Information (NMI), Adjusted Rand Index (ARI), and Accuracy (ACC). The backbone performance is compared across several configurations to highlight the individual and combined impact of the modules on feature extraction.
Table 8: Effect of cosine decay schedule (Sch.) and gradient clipping (Clip.) on feature extraction.
Dataset Scheduler & Backbone
Grad Clipping NMI ARI ACC
STL-10 - $\mathbf{0.677\pm 0.020}$ $\mathbf{0.547\pm 0.031}$ $\mathbf{0.705\pm 0.039}$
$0.674\pm 0.015$ $0.545\pm 0.032$ $0.701\pm 0.034$
ImageNet-10 - $0.857\pm 0.021$ $0.772\pm 0.068$ $0.858\pm 0.071$
$\mathbf{0.886\pm 0.008}$ $\mathbf{0.844\pm 0.022}$ $\mathbf{0.918\pm 0.026}$
Table 9: Effect of feature-level head on feature extraction.
Dataset Feature-Level Backbone
NMI ARI ACC
STL-10 - $0.659\pm 0.016$ $0.530\pm 0.028$ $0.692\pm 0.030$
$\mathbf{0.674\pm 0.015}$ $\mathbf{0.545\pm 0.032}$ $\mathbf{0.701\pm 0.034}$
ImageNet-10 - $0.870\pm 0.022$ $0.809\pm 0.064$ $0.886\pm 0.058$
$\mathbf{0.886\pm 0.008}$ $\mathbf{0.844\pm 0.022}$ $\mathbf{0.918\pm 0.026}$
Table 10: Effect of single-view vs. dual-view data augmentation on feature extraction.
Dataset Augmentation Backbone
x1 x2 NMI ARI ACC
STL-10 - $0.660\pm 0.011$ $0.520\pm 0.022$ $0.683\pm 0.022$
$\mathbf{0.671\pm 0.019}$ $\mathbf{0.545\pm 0.032}$ $\mathbf{0.701\pm 0.034}$
ImageNet-10 - $\mathbf{0.889\pm 0.010}$ $\mathbf{0.851\pm 0.024}$ $\mathbf{0.923\pm 0.018}$
$0.886\pm 0.008$ $0.844\pm 0.022$ $0.918\pm 0.026$

### v2

@@
-Contrastive Disentangling: Fine-grained representation learning through multi-level contrastive learning without class priors
+Fine-Grained Representation Learning via Multi-Level Contrastive Learning without Class Priors
 Abstract
-Recent advancements in unsupervised representation learning often leverage class information to enhance feature extraction and clustering performance.
-However, this reliance on class priors limits the applicability of such methods in real-world scenarios where class information is unavailable or ambiguous.
-In this paper, we propose Contrastive Disentangling (CD), a simple and effective framework that learns representations without any reliance on class priors.
-Our framework employs a multi-level contrastive learning strategy that combines instance-level and feature-level losses with a normalized entropy loss to learn semantically rich and fine-grained representations.
-Specifically, (1) the instance-level contrastive loss encourages the separation of feature representations for different samples, (2) the feature-level contrastive loss promotes independence among the feature head predictions, and (3) the normalized entropy loss encourages the feature heads to capture meaningful and prevalent attributes from the data.
-These components work together to enable CD to significantly outperform existing methods, as demonstrated by extensive experiments on benchmark datasets including CIFAR-10, CIFAR-100, STL-10, and ImageNet-10, particularly in scenarios where class priors are absent.
+Recent advances in unsupervised representation learning frequently leverage class information to improve the extraction and clustering of features.
+However, this dependence on class priors limits the applicability of such methods in real-world scenarios where class information is unavailable or ambiguous.
+In this paper, we propose Contrastive Disentangling (CD), a simple yet effective framework that learns representations without any relying on class priors.
+CD employs a multi-level contrastive learning strategy, integrating instance-level and feature-level losses with a normalized entropy loss to learn semantically rich and fine-grained representations.
+Specifically, (1) the instance-level contrastive loss encourages the separation of feature representations between different samples; (2) the feature-level contrastive loss promotes independence among feature prediction heads; and (3) the normalized entropy loss ensures that the feature heads capture meaningful and prevalent attributes from the data.
+These components together enable CD to outperform existing methods in scenarios lacking class priors, as demonstrated by extensive experiments on benchmark datasets including CIFAR-10, CIFAR-100, STL-10, and ImageNet-10.
 The code is available at https://github.com/Hoper-J/Contrastive-Disentangling.
 keywords:
-Disentangling representations , Clustering , Contrastive learning , Representation learning , Unsupervised learning
+Disentangled representations , Clustering , Contrastive learning , Representation learning , Unsupervised learning
 […]
@@
 1 Introduction
-The growing importance of unsupervised learning in processing large-scale unlabeled datasets has increased the demand for effective feature extraction techniques that do not rely on predefined class labels.
-In computer vision tasks, unsupervised representation learning has become a powerful solution for deriving meaningful feature representations from raw data.
-A key advancement in this domain is contrastive learning, which has emerged as a prominent paradigm within unsupervised representation learning.
-By maximizing consistency between different augmented views of the same data instance while minimizing similarity between different instances, contrastive learning has significantly advanced feature extraction (Ye et al., 2019; Chen et al., 2020).
-As research in this field progresses, numerous contrastive learning methods (Van Gansbeke et al., 2020; Li et al., 2021; Qian, 2023) have started to utilize class number information to further improve performance.
-These approaches use class numbers as prior knowledge during network construction, achieving state-of-the-art results in unsupervised clustering tasks.
-This suggests that incorporating class information during training yield more discriminative feature representations.
-However, in practical applications, class information is often unavailable, or the data may lack clear structure, limiting the applicability of these methods.
-To address this limitation, we propose a novel framework, Contrastive Disentangling (CD), which enhances feature extraction and disentanglement in unsupervised learning tasks through a multi-level contrastive learning strategy and an innovative disentanglement loss function, all without relying on class number information.
-Unlike methods that rely on additional complex modules, such as self-labeling, our approach focuses on validating the effectiveness of the core framework to ensure its broad applicability.
-Figure 1: Multi-level Contrastive Learning in the CD Framework. Instance-level (top) operates on image instances, while feature-level (bottom) focuses on features visualized with LIME.
-As illustrated in Figure 1, the CD framework incorporates two levels of contrastive learning: instance-level and feature-level.
-Instance-level contrastive learning aims to distinguish different image instances within the feature space, while feature-level contrastive learning seeks to differentiate the features captured by distinct prediction heads.
-In the feature-level module, we employ the LIME (Local Interpretable Model-Agnostic Explanations) technique (Ribeiro et al., 2016) to interpret the outputs of specific feature prediction heads, providing an intuitive visualization of the diverse feature attributes captured by the model across the image batch.
-In summary, the main contributions of this paper include:
+The growing availability of large-scale unlabeled datasets has underscored the importance of unsupervised learning in extracting meaningful features without predefined class labels.
+In computer vision, contrastive learning has become a dominant method for unsupervised representation learning, improving feature extraction by maximizing consistency between augmented views of the same instance while minimizing similarity between distinct instances [1, 2].
+Recent contrastive learning methods [3, 4, 5] have incorporated class priors to enhance performance.
+By using class counts as prior knowledge during network construction, these approaches achieve state-of-the-art results in unsupervised clustering tasks, indicating that class information can lead to more discriminative feature representations.
+However, this reliance on class priors limits their applicability in real-world scenarios where class information is uncertain.
+To address this limitation, we propose Contrastive Disentangling (CD), a novel framework that refines feature extraction and disentanglement in unsupervised learning without relying on class priors.
+CD integrates multi-level contrastive learning with a disentanglement loss to ensure feature diversity and independence, making it applicable to scenarios lacking class information.
+Our approach avoids complex additional modules such as pseudo-labeling, focusing instead on the effectiveness of the core framework.
+Figure 1: Overview of the Multi-level Contrastive Learning Strategy in CD Framework.
+The instance-level (top) operates on image instances, while the feature-level (bottom) focuses on features visualized using LIME.
+As shown in Figure 1, CD utilizes two levels of contrastive learning: instance-level and feature-level.
+Instance-level contrastive learning distinguishes image instances in the feature space, while feature-level contrastive learning ensures that distinct prediction heads capture diverse and complementary feature attributes.
+The feature-level module employs the LIME (Local Interpretable Model-Agnostic Explanations) technique [6] to visualize feature predictions, offering an intuitive interpretation of the model’s behavior.
+In summary, the main contributions of this paper are:
   1. 1.
-    We introduce feature prediction heads, a novel addition to the feature learning process.
-Using normalized entropy loss, these heads effectively capture common features prevalent in the dataset, thereby enhancing the representational capacity and generalization of the features.
+    Feature Prediction Heads: We introduce feature prediction heads as a novel component in the feature learning process.
+Through the use of a normalized entropy loss, these heads effectively capture common features prevalent in the dataset, enhancing the representational capacity and generalization of the learned features.
   2. 2.
-    A feature-level contrastive learning strategy is proposed, ensuring that different feature prediction heads focus on learning distinct and complementary features, thus promoting feature disentanglement.
+    Feature-Level Contrastive Learning: We propose a feature-level contrastive learning strategy to encourage different feature prediction heads to focus on distinct and complementary features, promoting effective feature disentanglement.
   3. 3.
-    We propose a simple and easy-to-use framework that significantly enhances the learning of fine-grained features in unsupervised learning tasks without relying on class information.
-Extensive experiments across multiple benchmark datasets demonstrate the superior performance of the CD framework, particularly in metrics like NMI, where it outperforms compared methods.
-Through comprehensive ablation studies, we validate the effectiveness of each component, providing new insights in unsupervised representation learning.
+    Simple and Effective Framework: We present a straightforward and effective framework that improves fine-grained feature learning in unsupervised tasks without relying on class priors.
+Extensive experiments on benchmark datasets demonstrate the superiority of CD, particularly in NMI scores.
+Ablation studies confirm the effectiveness of each component, providing new insights into unsupervised representation learning.
 2 Related Work
@@
 Unsupervised representation learning seeks to extract meaningful features from unlabeled data.
-The field is primarily dominated by two approaches: generative models and contrastive learning.
-Generative models (Bengio et al., 2006; Kingma, 2013; Goodfellow et al., 2014) learn to generate samples resembling the true data distribution by modeling the underlying latent space.
-However, these models often struggle to learn invariant feature representations, which are insensitive to certain input transformations like rotation, translation, and scaling (Ntelemis et al., 2023).
-To overcome these challenges, instance-level contrastive learning methods (He et al., 2020; Chen et al., 2020; Grill et al., 2020; Chen and He, 2021) have been proposed.
-These approaches significantly enhance the model’s capacity to extract invariant features by maximizing the similarity between different views of the same instance.
-These methods have exhibited superior performance across various downstream tasks, underscoring their efficacy in unsupervised learning scenarios.
-2.2 Utilization of Class Number Information
-Recent approaches have leveraged class number information as additional prior knowledge to enhance model performance.
-Methods (Van Gansbeke et al., 2020; Li et al., 2021; Niu et al., 2022; Qian, 2023) incorporate class number to constrain the final output during network construction.
-Several of these methods (Van Gansbeke et al., 2020; Niu et al., 2022; Qian, 2023) also employ pseudo-labeling techniques, further enhancing the model’s performance in unsupervised clustering tasks.
-These advancements indicate that class-level information can effectively assist models in distilling more semantically meaningful features.
-However, this dependence on class information constrains the applicability of these methods in real-world settings where class labels or the number of classes are frequently unknown or ambiguous.
-This limitation emphasizes the necessity for more adaptive approaches that can perform effectively without relying on class priors.
+This area is dominated by two main approaches: generative models and contrastive learning.
+Generative models [7, 8, 9] aim to generate samples that resemble the true data distribution by modeling the underlying latent space.
+Despite their success in data generation, these models often struggle to learn invariant feature representations—features that remain consistent under various input transformations such as rotation, translation, and scaling [10].
+To overcome these limitations, instance-level contrastive learning methods [2, 11, 12, 13] have been developed.
+These approaches enhance the model’s ability to extract invariant features by maximizing the similarity between different augmented views of the same instance while minimizing the similarity between different instances.
+As a result, contrastive learning has proven to be highly effective across various downstream tasks, offering robust performance in unsupervised learning.
+2.2 Utilization of Class Priors
+Several recent approaches [4, 5, 14] have incorporated class prior information to improve model performance by constraining the final outputs based on known class counts.
+These methods often achieve state-of-the-art results in unsupervised clustering tasks by guiding the model to learn more discriminative feature representations.
+Additionally, pseudo-labeling techniques are sometimes employed to iteratively refine cluster assignments, further boosting performance.
+However, these methods are limited by their reliance on class priors, making them unsuitable for real-world tasks where class labels are often unknown or ambiguous.
+This reliance highlights the need for methods that can operate effectively without class information.
 2.3 Multi-level Contrastive Learning
-The introduction of cluster-level contrastive learning by CC (Li et al., 2021) for clustering tasks has unveiled new avenues in contrastive learning research.
-However, partitioning based on predefined categories may not adequately capture the fine-grained features inherent in datasets, potentially constraining the model’s feature extraction capabilities.
-Building upon this observation, we propose a novel architecture that leverages feature space representations as inputs to feature prediction heads.
-By extracting common features and implementing feature-level contrastive learning, our approach achieves relative decoupling of distinct feature heads.
-This facilitates the capture of more discriminative and fine-grained features, thereby advancing the state-of-the-art in unsupervised representation learning.
+Multi-level contrastive learning has emerged as a promising direction within contrastive learning research.
+The introduction of cluster-level contrastive learning by Li et al.
+[4] for clustering tasks has opened new avenues for exploring more structured representations.
+However, methods that partition data based on predefined categories may miss the fine-grained features inherent in more complex datasets, limiting their ability to capture nuanced representations.
+Building on this insight, we propose the Contrastive Disentangling (CD) framework, which leverages feature space representations as inputs to multiple feature prediction heads.
+By combining feature-level and instance-level contrastive learning, CD encourages a decoupling of distinct feature heads, enabling the model to learn more diverse and fine-grained feature representations.
+This approach advances unsupervised representation learning without relying on class priors, offering a more flexible and robust solution for a variety of tasks.
 3 Method
 3.1 Proposed Contrastive Disentangling Framework
-The proposed CD framework leverages multi-level contrastive learning, as illustrated in Figure 2. It consists of the following key components:
+The proposed Contrastive Disentangling (CD) framework leverages a multi-level contrastive learning strategy to achieve fine-grained representation learning without the need for class priors, as illustrated in Figure 2.
 Figure 2: Architecture of the proposed Contrastive Disentangling (CD) framework.
-The top section illustrates the shared backbone network, instance-level projector, and feature-level predictor, while the bottom section displays the optimization objectives for instance-level and feature-level contrastive learning.
+The upper section illustrates the shared backbone network, instance-level projector, and feature predictor, while the lower section displays the optimization objectives for instance-level and feature-level contrastive learning.
+The CD framework comprises the following key components:
   • 1.
-    Data Augmentation.
-Each input batch $x$ undergoes a series of data augmentation operations, such as random cropping, horizontal flipping, color distortion, and blurring.
-These augmentations generate two correlated views, denoted as $x^{(1)}$ and $x^{(2)}$, enabling the model to focus on semantically relevant features that remain invariant under these transformations.
+    Data Augmentation: Each input batch $\mathbf{x}$ undergoes a series of data augmentation operations, including random cropping, horizontal flipping, color jittering, and Gaussian blurring.
+These augmentations generate two correlated views, denoted as $\mathbf{x}^{(1)}$ and $\mathbf{x}^{(2)}$, of the same sample.
+This dual-view approach encourages the model to focus on semantically relevant features that are invariant to such transformations.
   • 2.
-    Feature Extraction via ResNet Backbone.
-A ResNet architecture (He et al., 2016) is used as the shared backbone network $f(\cdot)$, which extracts feature representations from the augmented views $x^{(1)}$ and $x^{(2)}$.
-This choice aligns with prevailing practices in contrastive learning (Chen et al., 2020; Li et al., 2021).
-We denote the extracted feature representations as $h^{(1)}=f(x^{(1)})$ and $h^{(2)}=f(x^{(2)})$.
+    Feature Extraction: The augmented views are processed by a shared backbone network $f(\cdot)$, implemented using a ResNet architecture [15].
+This backbone extracts feature representations $\mathbf{h}^{(1)}=f(\mathbf{x}^{(1)})$ and $\mathbf{h}^{(2)}=f(\mathbf{x}^{(2)})$, ensuring consistent feature extraction across different views.
   • 3.
-    Instance Projector.
-The feature representations $h^{(1)}$ and $h^{(2)}$ are passed through an instance projector $g(\cdot)$, implemented as a multi-layer perceptron (MLP).
-This MLP consists of two linear layers, batch normalization (Ioffe, 2015), and a ReLU activation function.
-The output is a set of latent representations $z^{(1)}=g(h^{(1)})$ and $z^{(2)}=g(h^{(2)})$, which are crucial for the contrastive learning process (Chen et al., 2020).
+    Instance Projector: The extracted features are passed through an instance projector $g(\cdot)$, a multi-layer perceptron (MLP) consisting of two linear layers with batch normalization [16] and ReLU activation functions.
+This projector maps the features into a latent space, producing latent representations $\mathbf{z}^{(1)}=g(\mathbf{h}^{(1)})$ and $\mathbf{z}^{(2)}=g(\mathbf{h}^{(2)})$.
+These representations are then used for instance-level contrastive learning.
   • 4.
-    Feature Predictor.
-The latent representations $z^{(1)}$ and $z^{(2)}$ are fed into a feature predictor $h(\cdot)$, structured similarly to the instance projector but using sigmoid activation in the final layer.
-The feature predictor outputs predictions $y^{(1)}=h(z^{(1)})$ and $y^{(2)}=h(z^{(2)})$, organized into multiple prediction heads.
-Each head predicts a distinct feature attribute.
-We will validate the design of the feature predictor in Section 4.5.3.
+    Feature Predictor: To capture diverse and disentangled feature attributes, the latent representations are further processed by a feature predictor $h(\cdot)$.
+This module, structurally similar to the instance projector but with a sigmoid activation in the final layer, outputs predictions $\mathbf{y}^{(1)}=h(\mathbf{z}^{(1)})$ and $\mathbf{y}^{(2)}=h(\mathbf{z}^{(2)})$.
+Each $\mathbf{y}^{(v)}$ is a matrix of size $N\times K$, where $K$ denotes the number of feature heads.
+Each feature head is designed to capture a distinct attribute of the data.
   • 5.
-    Loss Functions.
-The total loss comprises three components: the instance-level contrastive loss, the feature-level contrastive loss, and an additional normalized entropy loss.
-These losses work together to enhance feature diversity, promote independence among learned features, and ensure a wide range of feature predictions.
+    Loss Functions: The total loss function of CD integrates three components: the instance-level contrastive loss $\mathcal{L}_{\text{inst}}$, the feature-level contrastive loss $\mathcal{L}_{\text{feat}}$, and the normalized entropy loss $\mathcal{L}_{\text{entropy}}$.
+These losses work synergistically to promote feature diversity, encourage independence among the learned features, and prevent feature collapse.
 3.2 Instance-level Contrastive Learning
-The objective of instance-level contrastive learning is to pull the representations of different augmented views of the same image closer in feature space while pushing apart representations of different images.
-Specifically, given an input batch $x$, two views $x^{(1)}$ and $x^{(2)}$ are generated via data augmentation, which are then processed by the shared backbone network and instance projector to yield latent representations $z^{(1)}=g(f(x^{(1)}))$ and $z^{(2)}=g(f(x^{(2)}))$.
-These latent representations, $z^{(1)}$ and $z^{(2)}$, are concatenated into a matrix $\mathbf{z}\in\mathbb{R}^{2N\times d}$, where $N$ is the batch size and $d$ is the feature dimension:
-$\mathbf{z}=\begin{bmatrix}z^{(1)}\\ z^{(2)}\end{bmatrix}=\begin{bmatrix}z_{1}\\ z_{2}\\ \vdots\\ z_{2N}\end{bmatrix}$ (1)
-To quantify the similarity between any two latent representations, we introduce the cosine similarity measure $\text{sim}_{z}(i,j)$:
-$\text{sim}_{z}(i,j)=\frac{z_{i}\cdot z_{j}}{\|z_{i}\|\|z_{j}\|}$ (2)
-where $i,j\in\{1,2,\dots,2N\}$ and $N$ is the batch size.
-Based on this similarity measure, we optimize the model by maximizing the similarity between different views of the same instance while minimizing the similarity between different instances.
-The loss function for instance $i$ is defined as:
+Instance-level contrastive learning aims to learn discriminative features by bringing representations of different augmented views of the same instance closer while pushing apart representations of different instances.
+This dual objective enhances the model’s ability to distinguish between distinct instances and recognize invariant features across different augmentations.
+Given an input batch $\mathbf{x}$ of size $N$, two augmented views $\mathbf{x}^{(1)}$ and $\mathbf{x}^{(2)}$ are generated.
+After passing through the backbone network and instance projector, latent representations $\mathbf{z}^{(1)}$ and $\mathbf{z}^{(2)}$ are obtained.
+These representations are concatenated to form a combined set of $2N$ latent vectors:
+$\mathbf{z}=\begin{bmatrix}\mathbf{z}^{(1)}\\ \mathbf{z}^{(2)}\end{bmatrix}=\begin{bmatrix}\mathbf{z}_{1}\\ \mathbf{z}_{2}\\ \vdots\\ \mathbf{z}_{2N}\end{bmatrix}\in\mathbb{R}^{2N\times d},$ (1)
+where $d$ is the dimensionality of the latent space.
+The similarity between any two latent representations is measured using cosine similarity:
+$\text{sim}_{z}(i,j)=\frac{\mathbf{z}_{i}\cdot\mathbf{z}_{j}}{\|\mathbf{z}_{i}\|\|\mathbf{z}_{j}\|}$ (2)
+for $i,j\in\{1,2,\dots,2N\}$.
+The purpose of using cosine similarity is to measure how closely related two feature vectors are, with a higher value indicating greater similarity.
+Building on this similarity measure, we define the instance-level contrastive loss $\mathcal{L}_{\text{inst}}$, which aims to maximize the similarity between positive pairs (i.e., the other augmented view of the same instance) and minimize the similarity between negative pairs (i.e., views of different instances).
+For each sample $i$, the instance-level contrastive loss is formulated as:
 $\ell_{i}^{\text{inst}}=-\log\frac{\exp(\text{sim}_{z}(i,\text{pos}(i))/\tau_{\text{inst}})}{\sum_{k=1}^{2N}\mathbbm{1}_{[k\neq i]}\exp(\text{sim}_{z}(i,k)/\tau_{\text{inst}})}$ (3)
-Here, $\text{pos}(i)$ represents the index of the positive sample corresponding to $i$.
-If $i\leq N$, then $\text{pos}(i)=i+N$; otherwise, $\text{pos}(i)=i-N$.
-The temperature parameter $\tau_{\text{inst}}$ within the instance-level contrastive loss governs the smoothness of the similarity distribution.
-Additionally, the indicator function $\mathbbm{1}_{[k\neq i]}$ assigns a value of 1 when $k\neq i$, thereby excluding the self-similarity of $z_{i}$.
-The instance-level contrastive loss is then calculated across all instance views:
+where $\text{pos}(i)$ denotes the index of the positive sample corresponding to $i$, $\tau_{\text{inst}}$ is a temperature parameter controlling the concentration level of the distribution, and $\mathbbm{1}_{[k\neq i]}$ is an indicator function that equals 1 when $k\neq i$ and 0 otherwise.
+This formulation follows the NT-Xent loss used in previous work [2, 17, 18, 19].
+The total instance-level contrastive loss is computed by averaging over all $2N$ samples:
 $\mathcal{L}_{\text{inst}}=\frac{1}{2N}\sum_{i=1}^{2N}\ell_{i}^{\text{inst}}$ (4)
 3.3 Disentanglement Loss
-To enhance the independence and diversity of the learned features, we introduce two complementary loss functions: feature-level contrastive loss and normalized entropy loss.
-These losses work together to promote the learning of disentangled and fine-grained feature attributes.
-Feature-level Contrastive Loss.
-The feature predictor $h(\cdot)$ is designed with multiple independent prediction heads, each focused on predicting a fine-grained feature attribute.
-After processing the input views $x^{(1)}$ and $x^{(2)}$ through the backbone network and instance projector to generate latent representations $z^{(1)}$ and $z^{(2)}$, these representations are fed into the feature predictor $h(\cdot)$, yielding the corresponding feature predictions $y^{(1)}$ and $y^{(2)}$.
-To align with the instance-level contrastive loss, these predictions $y$ are transposed and organized into a matrix $\mathbf{y}\in\mathbb{R}^{2K\times N}$, where $K$ denotes the number of features and $N$ represents the batch size:
-$\mathbf{y}=\begin{bmatrix}y^{(1)\top}\\ y^{(2)\top}\end{bmatrix}=\begin{bmatrix}y_{1}\\ y_{2}\\ \vdots\\ y_{2K}\end{bmatrix}$ (5)
-In this matrix, $y_{i}$ and $y_{\text{pos}(i)}$ correspond to the prediction results of feature head $i$ on the same batch of samples under two different views.
-We then define the similarity function $\text{sim}_{y}(i,j)$ to measure the cosine similarity between any two feature prediction vectors:
-$\text{sim}_{y}(i,j)=\frac{y_{i}\cdot y_{j}}{\|y_{i}\|\|y_{j}\|}$ (6)
-To ensure that feature head $i$ learns features different from other heads, the feature-level contrastive loss is defined as follows:
+To ensure that the learned feature representations are both diverse and independent, we introduce two complementary loss functions: the feature-level contrastive loss, which promotes diversity by encouraging each feature head to focus on distinct attributes, and the normalized entropy loss, which ensures that the captured features are meaningful and prevalent across the dataset.
+3.3.1 Feature-Level Contrastive Loss
+The feature predictor $h(\cdot)$ produces predictions $\mathbf{y}^{(1)}$ and $\mathbf{y}^{(2)}$, each of size $N\times K$, where $K$ is the number of feature heads.
+To apply feature-level contrastive learning, we concatenate the transposed outputs of the two views:
+$\mathbf{y}=\begin{bmatrix}(\mathbf{y}^{(1)})^{\top}\\ (\mathbf{y}^{(2)})^{\top}\end{bmatrix}=\begin{bmatrix}\mathbf{y}_{1}\\ \mathbf{y}_{2}\\ \vdots\\ \mathbf{y}_{2K}\end{bmatrix}\in\mathbb{R}^{2K\times N},$ (5)
+where $\mathbf{y}_{i}$ represents the predictions of feature head $i$ across the batch. The cosine similarity between any two feature prediction vectors is defined as:
+$\text{sim}_{y}(i,j)=\frac{\mathbf{y}_{i}\cdot\mathbf{y}_{j}}{\|\mathbf{y}_{i}\|\|\mathbf{y}_{j}\|}$ (6)
+The feature-level contrastive loss encourages each feature head to capture distinct attributes by maximizing the similarity between corresponding heads across different views while minimizing the similarity between different heads.
+For feature head $i$, the loss is defined as:
 $\ell_{i}^{\text{feat}}=-\log\frac{\exp(\text{sim}_{y}(i,\text{pos}(i))/\tau_{\text{feat}})}{\sum_{k=1}^{2K}\mathbbm{1}_{[k\neq i]}\exp(\text{sim}_{y}(i,k)/\tau_{\text{feat}})}$ (7)
-Here, $\tau_{\text{feat}}$ is the temperature parameter for the feature-level contrastive loss, which controls the degree of similarity enforced between different features.
-The primary objective is to encourage each prediction head to learn disentangled feature attributes.
-The final feature-level contrastive loss is calculated as follows:
+where $\text{pos}(i)$ denotes the index of the positive sample for feature head $i$, and $\tau_{\text{feat}}$ is the temperature parameter for the feature-level contrastive loss.
+The total feature-level contrastive loss is computed by averaging over all $2K$ feature heads:
 $\mathcal{L}_{\text{feat}}=\frac{1}{2K}\sum_{i=1}^{2K}\ell_{i}^{\text{feat}}$ (8)
-Normalized Entropy Loss.
-To further promote the learning of common features, we introduce a normalized entropy loss $\mathcal{L}_{\text{entropy}}$.
-This loss encourages the model to produce balanced and varied predictions by maximizing the entropy of the predicted outputs.
-By doing so, it prevents the model from overfitting to specific features and promotes the capture of more generalizable attributes.
-The normalized binary entropy $H(y_{i})$ for each feature prediction vector $y_{i}$ is defined as:
-$H(y_{i})=-\frac{1}{N\log(2)}\sum_{j=1}^{N}\left(y_{ij}\log(y_{ij})+(1-y_{ij})\log(1-y_{ij})\right)$ (9)
-Subsequently, the normalized entropy loss $\mathcal{L}_{\text{entropy}}$ is computed as the mean entropy across all feature predictions:
-$\mathcal{L}_{\text{entropy}}=\frac{1}{2K}\sum_{i=1}^{2K}H(y_{i})$ (10)
-Finally, the disentanglement loss $\mathcal{L}_{\text{disentanglement}}$ combines the feature-level contrastive loss and the normalized entropy loss:
-$\mathcal{L}_{\text{disentanglement}}=\mathcal{L}_{\text{feat}}-\alpha\cdot\mathcal{L}_{\text{entropy}}$ (11)
-where $\alpha$ is a hyperparameter that controls the weight of the normalized entropy loss.
-It balances the learning of common features within the dataset and the disentanglement of independent feature representations.
-Minimizing this combined loss enables our CD framework to effectively facilitate the unsupervised extraction and disentanglement of meaningful features from the input data.
-The efficacy of $\mathcal{L}_{\text{disentanglement}}$ is further validated in the experimental section 4.5.3.
+3.3.2 Normalized Entropy Loss
+To prevent feature collapse and ensure that the learned features are meaningful and widespread across the dataset, we introduce a normalized entropy loss $\mathcal{L}_{\text{entropy}}$.
+This loss encourages each feature head to distribute its predictions evenly, capturing non-trivial and prevalent patterns in the data.
+The normalized binary entropy for each feature prediction vector $\mathbf{y}_{i}$ is defined as:
+$H(\mathbf{y}_{i})=-\frac{1}{N\log(2)}\sum_{j=1}^{N}\left(\mathbf{y}_{ij}\log(\mathbf{y}_{ij})+(1-\mathbf{y}_{ij})\log(1-\mathbf{y}_{ij})\right)$ (9)
+The overall normalized entropy loss is the average entropy across all feature heads:
+$\mathcal{L}_{\text{entropy}}=\frac{1}{2K}\sum_{i=1}^{2K}H(\mathbf{y}_{i})$ (10)
+3.3.3 Combined Disentanglement Loss
+The total disentanglement loss is the combination of the feature-level contrastive loss and the normalized entropy loss:
+$\mathcal{L}_{\text{disentanglement}}=\mathcal{L}_{\text{feat}}-\alpha\mathcal{L}_{\text{entropy}}$ (11)
+where $\alpha$ is a hyperparameter that controls the balance between encouraging diversity and ensuring meaningful feature representations.
+By jointly optimizing these two objectives, the feature heads are encouraged to capture both diverse and informative patterns, leading to higher-quality feature disentanglement.
+3.4 Total Loss
+The total loss function for training the CD framework is the sum of the instance-level contrastive loss and the disentanglement loss.
+This ensures that both the instance-level discrimination and feature-level disentanglement are optimized simultaneously:
+$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{inst}}+\mathcal{L}_{\text{disentanglement}}.$ (12)
+By optimizing both the instance and feature-level objectives, our framework achieves a robust balance between learning discriminative instance-level representations and promoting diverse, meaningful feature representations without relying on class priors.
+Algorithm 1 summarizes the proposed method.
 Input: Dataset $X$, Batch size $N$, Feature number $K$, Temperature constants $\tau_{\text{inst}}$ and $\tau_{\text{feat}}$
@@
        end for
-      Combine $z^{(1)}$, $z^{(2)}$ by Eq.(1) and $y^{(1)}$, $y^{(2)}$ by Eq.(5);
+      Combine $\mathbf{z^{(1)}}$, $\mathbf{z^{(2)}}$ by Eq.(1) and $\mathbf{y^{(1)}}$, $\mathbf{y^{(2)}}$ by Eq.(5)
+      // Compute Losses
        Compute cosine similarities $\text{sim}_{z}(i,j)$, $\text{sim}_{y}(i,j)$ by Eq. (2), (6);
-      Calculate instance-level contrastive loss $\mathcal{L}_{\text{inst}}$ by Eq. (3), (4);
-      Calculate feature-level contrastive loss $\mathcal{L}_{\text{feat}}$ by Eq. (7), (8);
-      Calculate normalized entropy loss $\mathcal{L}_{\text{entropy}}$ by Eq. (9), (10);
-      Combine all losses $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{inst}}+\mathcal{L}_{\text{feat}}-\alpha\cdot\mathcal{L}_{\text{entropy}}$;
-      Update parameters of $f$, $g$, and $h$ to minimize $\mathcal{L}_{\text{total}}$;
+      Compute instance-level contrastive loss $\mathcal{L}_{\text{inst}}$ by Eq. (3), (4);
+      Compute feature-level contrastive loss $\mathcal{L}_{\text{feat}}$ by Eq. (7), (8);
+      Compute normalized entropy loss $\mathcal{L}_{\text{entropy}}$ by Eq. (9), (10);
+      Compute total loss $\mathcal{L}_{\text{total}}$ by Eq. (12);
+      // Update Model Parameters
+       Update $f$, $g$, and $h$ by minimizing $\mathcal{L}_{\text{total}}$;
 end for
@@
 4.1 Datasets
-We conducted our experiments on four widely-used computer vision datasets: CIFAR-10, CIFAR-100 (Krizhevsky and Hinton, 2009), STL-10 (Coates et al., 2011), and ImageNet-10 (Deng et al., 2009).
+We conducted experiments on four widely-used datasets: CIFAR-10, CIFAR-100 [20], STL-10 [21], and ImageNet-10 [22].
 These datasets serve as benchmarks to rigorously evaluate the effectiveness of our proposed Contrastive Disentangling (CD) framework.
-Table 1 provides an overview of the dataset splits, sample sizes, and the number of classes used.
+Table 1 provides an overview of the dataset splits, sample sizes, and the number of classes.
 Table 1: Overview of datasets used for training and evaluation
@@
 ImageNet-10 train 13,000 10
-For the CIFAR-100 dataset, we used 20 super-classes as ground-truth labels instead of the 100 fine-grained classes, aligning with benchmark settings in previous studies.
-Furthermore, to ensure the rigor of our architecture evaluation, we did not use any unseen data during training, such as the unlabeled data in STL-10.
+For CIFAR-100, we used 20 super-classes as ground-truth labels instead of the 100 fine-grained classes, aligning with benchmark settings in previous studies.
+Additionally, to guarantee the rigor of our assessment, no unseen data (e.g., the unlabeled data in STL-10) was used during training.
 4.2 Experimental Settings
 Data Augmentation.
-To ensure fair comparison with existing methods (Li et al., 2021), we adopted the same data augmentation strategies and used ResNet-34 as the backbone network for feature extraction.
-Specifically, all datasets underwent random cropping to 224x224, random horizontal flipping, color jittering, and grayscale transformation.
-Notably, Gaussian blur was applied only to the ImageNet-10 dataset, as upscaling images to 224x224 for smaller datasets (e.g., CIFAR-10, CIFAR-100, and STL-10) inherently introduces a similar blurring effect.
+To ensure comparability with existing methods [4], we adopted the same data augmentation strategies, using ResNet-34 as the backbone for feature extraction.
+All datasets were subject to random cropping to 224x224, horizontal flipping, color jittering, and grayscale transformation.
+Additionally, Gaussian blur was applied to the ImageNet-10 dataset, while the upscaling of smaller datasets (CIFAR-10, CIFAR-100, and STL-10) provided a comparable blurring effect.
 Training Details.
-In our experiments, the hidden dimension and the number of features were both set equal to the batch size.
-We employed a cosine decay schedule without restarts (Chen et al., 2020; Loshchilov and Hutter, 2016) and gradient clipping to enhance the stability of model training.
-The effectiveness of these optimizations is demonstrated in the ablation study presented in Table 2.
-Additionally, no other techniques (e.g., self-labeling) were used to boost model performance.
-The hyperparameter $\alpha$, which balances the diversity of feature representations, was simply set to 1.
-The remaining training settings were consistent with prior contrastive learning work (Li et al., 2021): a random seed of 42 was used, and training was conducted for 1000 epochs.
-The Adam optimizer (Kingma, 2014) was employed, with a learning rate of $3\times 10^{-4}$ and no weight decay.
-The temperature parameters $\tau_{\text{inst}}$ and $\tau_{\text{feat}}$ were set to 0.5 and 1, respectively.
+In our experiments, both the hidden dimension and the number of features were set equal to the batch size.
+A cosine decay schedule without restarts [2, 23] and gradient clipping were employed to stabilize training, as validated in our ablation studies (Table 5).
+No additional techniques, such as self-labeling, were used to improve performance.
+The temperature parameters $\tau_{\text{inst}}$ and $\tau_{\text{feat}}$ were set to 0.5 and 1, respectively, while the balance parameter $\alpha$ for feature diversity was set to 1.
+Other training settings were based on prior contrastive learning works, including 1000 training epochs, a random seed of 42, and the Adam optimizer [24] with a learning rate of $3\times 10^{-4}$ and no weight decay.
 Environment.
-The experiments were conducted on an NVIDIA RTX 3090 GPU (24GB) and an 18 vCPU AMD EPYC 9754 128-Core Processor.
-We experimented with batch sizes of 128 and 256.
-The training time for the CIFAR-10 and CIFAR-100 datasets was approximately 48 GPU hours, while for the STL-10 and ImageNet-10 datasets, the training time was around 11 GPU hours.
-To ensure reproducibility and transparency, detailed logs of all experimental procedures are available at here.
+The experiments were performed on an NVIDIA RTX 3090 GPU (24GB) and an 18 vCPU AMD EPYC 9754 128-Core Processor.
+Batch sizes of 128 and 256 were used.
+Training time for CIFAR-10 and CIFAR-100 was approximately 48 GPU hours, while for STL-10 and ImageNet-10, it was around 11 GPU hours.
+Detailed logs of all experimental procedures are available here.
 Evaluation Metrics.
-We employed three clustering metrics — Normalized Mutual Information (NMI), Adjusted Rand Index (ARI), and Accuracy (ACC) — to evaluate the feature extraction and disentanglement capabilities of the model.
+We used three clustering metrics — Normalized Mutual Information (NMI), Adjusted Rand Index (ARI), and Accuracy (ACC) — to evaluate the feature extraction and disentanglement capabilities of the model.
 4.3 Compared Methods
-To rigorously assess the performance of our proposed method in various unsupervised learning scenarios, we compared it against a diverse set of representative methods, chosen for their relevance to the challenges addressed by our approach.
-Alongside the traditional K-means (MacQueen et al., 1967) clustering algorithm, we evaluated several fully unsupervised models—AE (Bengio et al., 2006), DeCNN (Zeiler et al., 2010), DAE (Vincent et al., 2010), VAE (Kingma, 2013), and DCGAN (Radford, 2015)—that span a spectrum of approaches from classical techniques to modern deep generative models.
-These methods, like our proposed Contrastive Disentangling (CD) model, do not rely on any prior class information during training, including the number of classes.
-We also compared our method with unsupervised models that incorporate class number information—such as JULE (Yang et al., 2016), DEC (Xie et al., 2016), DAC (Chang et al., 2017), DCCM (Wu et al., 2019), PICA (Huang et al., 2020), and CC (Li et al., 2021)—to illustrate how our approach performs relative to methods that leverage this additional knowledge.
-To ensure a fair and consistent comparison, all experiments were conducted using the full set of samples, adhering to identical training and evaluation protocols.
+We compared CD against a diverse set of representative unsupervised models, selected based on their relevance to the challenges addressed by our approach.
+In addition to the traditional K-means clustering algorithm [25], we assessed fully unsupervised models such as AE [7], DeCNN [26], DAE [27], VAE [8], and DCGAN [28], all of which do not rely on class information during training, like our CD model.
+Furthermore, we benchmarked CD against unsupervised models that incorporate class number information, including JULE [29], DEC [30], DAC [31], DCCM [32], PICA [33], and CC [4].
+All experiments were conducted using the full dataset and adhered to identical training and evaluation protocols to ensure fairness.
 4.4 Experimental Results
-We evaluated our model at two key stages: feature extraction and feature disentanglement.
-Specifically, we applied K-means clustering to both the backbone network output and the final model output to assess the model’s capabilities in feature extraction and disentanglement, respectively.
-Feature Extraction.
-Table 2 presents a comparative analysis of various models at the feature extraction stage, utilizing the NMI, ARI, and ACC metrics.
-Across four datasets—CIFAR-10, CIFAR-100-20, STL-10, and ImageNet-10—our Contrastive Disentanglement (CD) method, implemented with batch sizes of 128 (CD-128) and 256 (CD-256), demonstrated superior performance compared to other fully unsupervised learning models, including AE, DeCNN, DAE, VAE, and DCGAN.
-In comparison to the contrastive learning method CC, which utilizes prior information about the number of classes, CD consistently produced the highest NMI scores across all datasets, with notable improvements in STL-10 and ImageNet-10, while maintaining superior overall performance in most other cases.
-Table 2: Feature extraction performance comparison across models. The 1st/2nd best results are indicated in bold/underlined
-Dataset Metrics K-means AE DeCNN DAE VAE DCGAN CC CD-128 CD-256
+We evaluated CD at two key stages: feature extraction and feature disentanglement. K-means clustering was applied to both the backbone output and the final model output to assess performance.
+Backbone.
+Table 2 presents a comparative analysis at the feature extraction stage using NMI, ARI, and ACC metrics.
+Across all datasets — CIFAR-10, CIFAR-100, STL-10, and ImageNet-10 — both variants of CD (CD-128 and CD-256) outperformed fully unsupervised methods such as AE, DeCNN, DAE, VAE, and DCGAN.
+Notably, CD achieved the highest NMI scores on all datasets, with significant improvements on STL-10 and ImageNet-10, and superior overall performance compared to the class-dependent model CC.
+Table 2: Feature extraction performance comparison across models.
+The 1st/2nd best results are indicated in bold/underlined.
+Models that utilized class number information during training are marked with *.
+Dataset Metrics K-means AE DeCNN DAE VAE DCGAN CC* CD-128 CD-256
 […]
@@
 […]
-Feature Disentanglement.
-Table 3 presents a comparison of our method with other unsupervised methods that utilize number of classes information at the final output stage.
-The results demonstrate that the CD method exhibited excellent performance across all metrics, particularly on the STL-10 and ImageNet-10 datasets.
-For CIFAR-10 and CIFAR-100-20, although the CD method’s ARI and ACC metrics were slightly lower than those of the CC method, this may be attributed to the CD method not utilizing the number of classes information to constrain the final feature prediction.
-Notably, on the STL-10 dataset, while the PICA method achieved a slightly higher ACC (0.713) due to the use of additional unlabeled data, the CD-128 model reached a comparable ACC of 0.702 without using extra data, and the CD-256 model further improved this to 0.734.
+Final output.
+Table 3 compares CD to models that use class number information.
+CD demonstrated strong performance across all metrics, particularly on more complex datasets like STL-10 and ImageNet-10.
+While ARI and ACC were slightly lower on CIFAR-10 and CIFAR-100 compared to CC, this is expected as CD does not use class number information.
+Importantly, CD achieved an ACC of 0.734 with a batch size of 256, surpassing PICA (0.713) despite not using additional unlabeled data.
 Table 3: Final output performance comparison across models.
-The 1st/2nd best results are indicated in bold/underlined, and models that utilized unlabeled datasets during training for STL-10 are marked with
-Dataset Metrics JULE DEC DAC DCCM PICA CC CD-128 CD-256
+The 1st/2nd best results are indicated in bold/underlined.
+Models that utilized class number information during training are marked with *, while those that utilized unlabeled datasets for STL-10 are marked with .
+Dataset Metrics JULE* DEC* DAC* DCCM* PICA* CC* CD-128 CD-256
 […]
@@
 […]
-These results provide strong evidence that the CD method can achieve excellent performance without requiring the number of classes information.
-It is particularly noteworthy that the CD method exhibited outstanding performance even with smaller batch sizes.
-For a detailed visualization of the CD-256 model’s feature representations using t-SNE, please see A.
+These results strongly support the effectiveness of CD in achieving high performance without relying on class priors.
+Furthermore, CD demonstrated excellent performance even with smaller batch sizes.
+For a t-SNE visualization of the CD-256 model’s feature representations, refer to A.
 4.5 Ablation Studies
-To demonstrate the effectiveness of each component in our model, we conducted a series of ablation studies using the STL-10 and ImageNet-10 datasets with a batch size of 128.
-In sections 4.5.1 and 4.5.2, we incrementally added modules to validate their effectiveness based on the framework and objective function shown in Figure 2.
-In sections 4.5.3 and 4.5.4, we removed certain modules to verify their necessity.
+We conducted ablation studies on STL-10 and ImageNet-10 with a batch size of 128 to assess the contributions of different components of our model.
+In Sections 4.5.1 and 4.5.2, we incrementally added modules to validate their effectiveness, while in Sections 4.5.3 and 4.5.4, we removed certain modules to verify their necessity.
 4.5.1 Normalized Entropy Loss and Batch Normalization
-The impact of normalized entropy loss (NE) and batch normalization (BN) on model performance was initially evaluated, as shown in Table 4.
-The experimental results demonstrate that both normalized entropy loss and batch normalization significantly enhance model performance, with their synergistic application yielding even more substantial performance improvements.
-Notably, NE exhibited particularly significant improvements during the feature disentanglement phase, underscoring its pivotal role in facilitating effective feature disentanglement.
+Table 4 demonstrates the impact of the normalized entropy loss (NE) and batch normalization (BN) on performance.
+Both components significantly improved results, with further enhancements observed when combined.
 Table 4: Effect of normalized entropy loss (NE) and batch normalization (BN) on performance across backbone and final output.
@@
 $\mathbf{0.857\pm 0.021}$ $\mathbf{0.772\pm 0.068}$ $\mathbf{0.858\pm 0.071}$ $\mathbf{0.890\pm 0.004}$ $\mathbf{0.864\pm 0.006}$ $\mathbf{0.929\pm 0.004}$
-To further illustrate the impact of normalized entropy loss on feature disentanglement, a visual representation of the model outputs is presented in Figure 3.
-(a) Default
+To illustrate the effect of NE loss on feature disentanglement, Figure 3 visualizes the model’s outputs with and without NE loss.
+This visualization highlights the importance of NE loss in guiding the model to capture common features across the dataset.
+(a) NE loss
 (b) no NE loss
 Figure 3: Visualization of transposed model outputs with/without normalized entropy loss (NE loss).
-Each row corresponds to the prediction of a specific feature attribute across the entire batch.
-The left visualization (a) shows the output with NE loss enabled, focusing on common features, while the right visualization (b) shows the output without NE loss, capturing more unique, image-specific features.
-This visualization clearly demonstrates the effectiveness of NE loss in guiding the model to capture common features across the dataset, rather than focusing on unique features specific to individual images.
-The rows in Figure 3 represent the feature attributes learned by our model, as depicted in the feature-level part of Figure 1 and further visualized in B.
-This visual evidence corroborates the quantitative results and emphasizes the significance of NE loss in the proposed methodology.
+The left figure (a) shows outputs with NE loss, emphasizing common features, while the right figure (b) shows outputs without NE loss, capturing more image-specific features.
+The rows in Figure 3 represent the feature attributes learned by our model, as depicted in the feature-level part of Figure 1. Further visualization of these attributes using LIME is available in B.
 4.5.2 Training Tricks
-Following the validation of core modules, we further evaluated the contribution of cosine decay schedule and gradient clipping training techniques, as shown in Table 5.
+Next, we evaluated the contribution of cosine decay scheduling and gradient clipping (Table 5). The results show that both techniques enhanced performance and stabilized training.
 Table 5: Effect of cosine decay schedule (Sch.) and gradient clipping (Clip.) on model performance.
@@
 $\mathbf{0.898\pm 0.003}$ $\mathbf{0.868\pm 0.006}$ $\mathbf{0.931\pm 0.003}$
-The results show that both the cosine decay schedule and gradient clipping further enhance the model’s final performance and improve the stability of the results.
 4.5.3 Feature Prediction Head
-Next, we evaluated the effect of removing the feature prediction head, which directly reduces the model to an instance-level contrastive learning architecture, similar to SimCLR.
-This experiment was designed to verify the critical role of the disentanglement loss $\mathcal{L}_{\text{disentanglement}}$in the feature disentanglement process.
+We evaluated the effect of removing the feature prediction head, which reduces the model to an instance-level contrastive learning framework similar to SimCLR.
+As shown in Table 6, the feature prediction head significantly boosts performance, and its removal results in substantial degradation, highlighting the critical role of feature-level contrastive learning in effective disentanglement (see Table 9 for backbone results).
+In comparison to SimCLR under identical training settings, our model demonstrates significant improvements: on the STL-10 dataset, NMI increases by 13.4%, ARI by 23.1%, and ACC by 10.9%; on ImageNet-10, NMI improves by 26.8%, ARI by 110.2%, and ACC by 53.1%.
+These results present that our feature-level contrastive learning approach enables the model to capture more distinct and diverse semantic features, leading to improved clustering and representation quality in the final output.
 Table 6: Effect of feature-level head on model performance.
@@
 $\mathbf{0.898\pm 0.003}$ $\mathbf{0.868\pm 0.006}$ $\mathbf{0.931\pm 0.003}$
-As shown in Table 6, the results demonstrate that the feature prediction head significantly improves the model’s performance in the final layers (see Table 9 for backbone results).
-Removing this head leads to a noticeable degradation in performance.
-This confirms the critical contribution of the feature-level contrastive learning strategy, which enhances the model’s ability to disentangle features across different prediction heads.
 4.5.4 Data Augmentation
-This section verifies the necessity of the dual-view data augmentation strategy for feature disentanglement.
+Finally, we examined the necessity of the dual-view data augmentation strategy (Table 7).
+The results indicate that dual-view augmentation is critical for feature disentanglement, whereas single-view augmentation can still achieve satisfactory performance during feature extraction (see Table 10 for details).
 Table 7: Effect of single-view vs. dual-view data augmentation on model performance.
@@
 $\mathbf{0.898\pm 0.003}$ $\mathbf{0.868\pm 0.006}$ $\mathbf{0.931\pm 0.003}$
-As shown in Table 7, the results demonstrate that dual-view augmentation is crucial in the feature disentanglement stage for capturing common features in the dataset.
-However, in the feature extraction stage, single-view augmentation can achieve satisfactory performance (see Table 10 for details).
 5 Conclusion
-In this work, we proposed the Contrastive Disentangling (CD) framework, a novel approach for enhancing feature extraction and disentanglement in unsupervised learning without relying on prior knowledge of class numbers.
-The CD framework integrates multi-level contrastive learning with an innovative disentanglement loss function, emphasizing simplicity and core effectiveness.
-Experimental results demonstrate that CD consistently outperforms class-dependent methods with simple architectures (i.e., without additional structures such as self-labeling), particularly in NMI, across multiple benchmark datasets including CIFAR-10, CIFAR-100, STL-10, and ImageNet-10.
-Moreover, CD exhibits robust performance in small-batch settings, effectively capturing common features across diverse data distributions.
-Comprehensive ablation studies corroborate the efficacy of each framework component, validating its overall effectiveness and potential for wide-scale application in unsupervised learning scenarios.
+In this work, we propose a novel framework that eliminates the reliance on class number assumptions, enabling the model to operate without predefined task constraints.
+By integrating multi-level contrastive learning with a disentanglement loss, our framework captures more fine-grained features from the data.
+Our experiments reveal that CD achieves comparable performance to class-dependent methods on smaller datasets (CIFAR-10, CIFAR-100), and outperforms them on more complex datasets (STL-10, ImageNet-10), despite its simple architecture.
+Ablation studies further confirm the effectiveness of each component.
+Moreover, our framework ensures that the final model outputs are enriched with semantic information, enhancing the interpretability of the learned features.
 6 Acknowledgments
@@
 Appendix A t-SNE Visualization of Backbone and Feature Predictor Outputs
-In this appendix, we present the t-SNE (t-Distributed Stochastic Neighbor Embedding) visualizations (Van der Maaten and Hinton, 2008) of the outputs both the backbone feature extraction layer and the feature predictor of our model, across all datasets: CIFAR-10, CIFAR-100, STL-10, and ImageNet-10.
+In this appendix, we present the t-SNE (t-Distributed Stochastic Neighbor Embedding) visualizations [34] of the outputs both the backbone feature extraction layer and the feature predictor of our model, across all datasets: CIFAR-10, CIFAR-100, STL-10, and ImageNet-10.
 These visualizations are shown in Figures 4-7, where (a) represents the backbone output and (b) represents the feature predictor output for each dataset.
@@
 Appendix B Visualizing Feature Prediction Heads with LIME
-We utilize the LIME (Local Interpretable Model-Agnostic Explanations) technique (Ribeiro et al., 2016) to visualize the outputs of the feature prediction heads.
+We utilize the LIME (Local Interpretable Model-Agnostic Explanations) technique [6] to visualize the outputs of the feature prediction heads.
 LIME is configured with $num\_samples$=5000, meaning that for each image, LIME generates 5000 perturbed samples to create an interpretable explanation.
@@
 $0.886\pm 0.008$ $0.844\pm 0.022$ $0.918\pm 0.026$
+It can be observed that when evaluating the model’s ability as a pretext model, or more specifically, the feature extraction capability of the backbone, the scheduler, gradient clipping, and dual-view data augmentation are not essential.
+However, if we aim for the final output of the model to be meaningful, rather than focusing solely on intermediate states, these components become crucial.
+You may refer to Table 5 - Table 7, especially Table 6, where our model demonstrates significantly higher semantic expressiveness in the final layer.

### v3

@@
 Abstract
-Recent advances in unsupervised representation learning frequently leverage class information to improve the extraction and clustering of features.
-However, this dependence on class priors limits the applicability of such methods in real-world scenarios where class information is unavailable or ambiguous.
-In this paper, we propose Contrastive Disentangling (CD), a simple yet effective framework that learns representations without any relying on class priors.
-CD employs a multi-level contrastive learning strategy, integrating instance-level and feature-level losses with a normalized entropy loss to learn semantically rich and fine-grained representations.
-Specifically, (1) the instance-level contrastive loss encourages the separation of feature representations between different samples; (2) the feature-level contrastive loss promotes independence among feature prediction heads; and (3) the normalized entropy loss ensures that the feature heads capture meaningful and prevalent attributes from the data.
-These components together enable CD to outperform existing methods in scenarios lacking class priors, as demonstrated by extensive experiments on benchmark datasets including CIFAR-10, CIFAR-100, STL-10, and ImageNet-10.
+Recent advances in unsupervised representation learning often rely on knowing the number of classes to improve feature extraction and clustering.
+However, this assumption raises an important question: is the number of classes always necessary, and do class labels fully capture the fine-grained features within the data?
+In this paper, we propose Contrastive Disentangling (CD), a framework designed to learn representations without relying on class priors.
+CD leverages a multi-level contrastive learning strategy, integrating instance-level and feature-level contrastive losses with a normalized entropy loss to capture semantically rich and fine-grained representations.
+Specifically, (1) the instance-level contrastive loss separates feature representations across samples; (2) the feature-level contrastive loss promotes independence among feature heads; and (3) the normalized entropy loss ensures feature diversity and prevents feature collapse.
+Extensive experiments on CIFAR-10, CIFAR-100, STL-10, and ImageNet-10 demonstrate that CD outperforms existing methods in scenarios where class information is unavailable or ambiguous.
 The code is available at https://github.com/Hoper-J/Contrastive-Disentangling.
@@
 In computer vision, contrastive learning has become a dominant method for unsupervised representation learning, improving feature extraction by maximizing consistency between augmented views of the same instance while minimizing similarity between distinct instances [1, 2].
-Recent contrastive learning methods [3, 4, 5] have incorporated class priors to enhance performance.
-By using class counts as prior knowledge during network construction, these approaches achieve state-of-the-art results in unsupervised clustering tasks, indicating that class information can lead to more discriminative feature representations.
-However, this reliance on class priors limits their applicability in real-world scenarios where class information is uncertain.
-To address this limitation, we propose Contrastive Disentangling (CD), a novel framework that refines feature extraction and disentanglement in unsupervised learning without relying on class priors.
-CD integrates multi-level contrastive learning with a disentanglement loss to ensure feature diversity and independence, making it applicable to scenarios lacking class information.
-Our approach avoids complex additional modules such as pseudo-labeling, focusing instead on the effectiveness of the core framework.
+Although recent contrastive learning methods [3, 4, 5] have leveraged class priors to improve performance, their reliance on predefined class counts limits their applicability in real-world scenarios where class information may be uncertain or unavailable.
+This dependence not only restricts the adaptability of these models but also hampers their ability to generalize to diverse tasks that lack explicit class labels.
+To address this limitation, we propose Contrastive Disentangling (CD), a novel unsupervised learning framework that improves feature extraction and disentanglement without requiring class priors.
 Figure 1: Overview of the Multi-level Contrastive Learning Strategy in CD Framework.
 The instance-level (top) operates on image instances, while the feature-level (bottom) focuses on features visualized using LIME.
-As shown in Figure 1, CD utilizes two levels of contrastive learning: instance-level and feature-level.
-Instance-level contrastive learning distinguishes image instances in the feature space, while feature-level contrastive learning ensures that distinct prediction heads capture diverse and complementary feature attributes.
-The feature-level module employs the LIME (Local Interpretable Model-Agnostic Explanations) technique [6] to visualize feature predictions, offering an intuitive interpretation of the model’s behavior.
+Our approach stems from a detailed observation of the model’s outputs.
+As illustrated in Figure 1, the rows of the instance projector correspond to the latent representations of the samples, while the columns of the feature predictor represent the predictions of fine-grained features across the entire batch.
+Traditional contrastive learning methods focus predominantly on instance-level feature representations, yet we raise a new question: can the focus be shifted to the features themselves?
+Motivated by this insight, we extend contrastive learning to the feature level, treating each feature prediction head as an independent “feature instance” to better capture the underlying attributes of the data.
+This conceptual shift forms the basis of our CD framework, which integrates multi-level contrastive learning with disentanglement loss to ensure feature diversity and independence, while maintaining a simple and efficient architecture.
 In summary, the main contributions of this paper are:
   1. 1.
-    Feature Prediction Heads: We introduce feature prediction heads as a novel component in the feature learning process.
-Through the use of a normalized entropy loss, these heads effectively capture common features prevalent in the dataset, enhancing the representational capacity and generalization of the learned features.
+    We introduce feature prediction heads and a novel feature-level contrastive approach, enabling the model to capture finer-grained features within a simple and effective framework, providing a fresh and generalizable perspective on contrastive learning.
   2. 2.
-    Feature-Level Contrastive Learning: We propose a feature-level contrastive learning strategy to encourage different feature prediction heads to focus on distinct and complementary features, promoting effective feature disentanglement.
+    We incorporate a normalized entropy loss to complement feature-level contrastive learning, ensuring that the model captures prevalent and diverse features across the dataset.
   3. 3.
-    Simple and Effective Framework: We present a straightforward and effective framework that improves fine-grained feature learning in unsupervised tasks without relying on class priors.
-Extensive experiments on benchmark datasets demonstrate the superiority of CD, particularly in NMI scores.
-Ablation studies confirm the effectiveness of each component, providing new insights into unsupervised representation learning.
+    To the best of our knowledge, this is the first contrastive learning method that achieves semantically rich final outputs without relying on class priors.
+By defining the model architecture independently of class information, our approach allows researchers to intuitively identify images with similar fine-grained features by examining the final model outputs, which exhibit the same level of semantic richness as the feature extraction layer.
+Moreover, the model offers researchers the flexibility to adjust the number of feature heads, enabling control over the granularity of the features learned.
+Our method demonstrates superior feature extraction capabilities compared to existing unsupervised frameworks and remains highly competitive against methods that rely on class information to constrain model outputs.
+Additionally, under identical training settings, the final outputs of CD exhibit richer semantics than SimCLR.
+This improvement is particularly evident in clustering tasks.
+For example, in the STL-10 dataset, CD achieves a 13.4% increase in NMI, a 23.1% increase in ARI, and a 10.9% improvement in ACC.
+On the ImageNet-10 dataset, CD shows even greater enhancements, with a 26.8% increase in NMI, a 110.2% improvement in ARI, and a 53.1% increase in ACC.
 2 Related Work
 2.1 Unsupervised Representation Learning
-Unsupervised representation learning seeks to extract meaningful features from unlabeled data.
-This area is dominated by two main approaches: generative models and contrastive learning.
+Unsupervised representation learning focuses on extracting meaningful features from unlabeled data, with two primary approaches: generative models and contrastive learning.
 Generative models [6, 7, 8] aim to generate samples that resemble the true data distribution by modeling the underlying latent space.
 Despite their success in data generation, these models often struggle to learn invariant feature representations—features that remain consistent under various input transformations such as rotation, translation, and scaling [9].
-To overcome these limitations, instance-level contrastive learning methods [2, 11, 12, 13] have been developed.
-These approaches enhance the model’s ability to extract invariant features by maximizing the similarity between different augmented views of the same instance while minimizing the similarity between different instances.
-As a result, contrastive learning has proven to be highly effective across various downstream tasks, offering robust performance in unsupervised learning.
+In contrast, instance-level contrastive learning methods [2, 10, 11, 12] improve the model’s ability to extract invariant features by maximizing the similarity between augmented views of the same instance while minimizing the similarity between different instances.
+This strategy has consistently demonstrated strong performance across various downstream tasks, positioning contrastive learning as a highly effective method in unsupervised learning.
 2.2 Utilization of Class Priors
-Several recent approaches [4, 5, 14] have incorporated class prior information to improve model performance by constraining the final outputs based on known class counts.
-These methods often achieve state-of-the-art results in unsupervised clustering tasks by guiding the model to learn more discriminative feature representations.
-Additionally, pseudo-labeling techniques are sometimes employed to iteratively refine cluster assignments, further boosting performance.
-However, these methods are limited by their reliance on class priors, making them unsuitable for real-world tasks where class labels are often unknown or ambiguous.
-This reliance highlights the need for methods that can operate effectively without class information.
+Recent methods [4, 5, 13] incorporate class priors to improve performance by constraining the learning process with known class counts and refining cluster assignments through pseudo-labeling.
+These approaches have yielded state-of-the-art results in unsupervised clustering, as class priors guide the model towards more discriminative representations.
+However, such reliance on class priors significantly limits the adaptability of these models, especially in real-world scenarios where class information is incomplete, ambiguous, or unavailable.
+This motivates the need for approaches that can effectively capture meaningful and discriminative representations without relying on predefined class information.
 2.3 Multi-level Contrastive Learning
-Multi-level contrastive learning has emerged as a promising direction within contrastive learning research.
-The introduction of cluster-level contrastive learning by Li et al.
-[4] for clustering tasks has opened new avenues for exploring more structured representations.
-However, methods that partition data based on predefined categories may miss the fine-grained features inherent in more complex datasets, limiting their ability to capture nuanced representations.
-Building on this insight, we propose the Contrastive Disentangling (CD) framework, which leverages feature space representations as inputs to multiple feature prediction heads.
-By combining feature-level and instance-level contrastive learning, CD encourages a decoupling of distinct feature heads, enabling the model to learn more diverse and fine-grained feature representations.
-This approach advances unsupervised representation learning without relying on class priors, offering a more flexible and robust solution for a variety of tasks.
+Multi-level contrastive learning has recently emerged as a novel direction within contrastive learning research.
+For example, cluster-level contrastive learning by Li et al.
+[4] was introduced for clustering tasks, opening new avenues for structured representation learning.
+However, methods that partition data based on predefined categories may overlook fine-grained features in more complex datasets, limiting their ability to capture subtle and nuanced patterns.
+Building on these observations, our framework extends contrastive learning to both the instance and feature levels.
+By treating each feature prediction head as an independent “feature instance,” our approach encourages each head to capture distinct, complementary attributes, which is a significant departure from traditional instance-level contrastive learning.
+This decoupling of feature heads allows CD to learn diverse and fine-grained representations, making it a flexible and powerful method for unsupervised representation learning without relying on class priors.
 3 Method
 3.1 Proposed Contrastive Disentangling Framework
-The proposed Contrastive Disentangling (CD) framework leverages a multi-level contrastive learning strategy to achieve fine-grained representation learning without the need for class priors, as illustrated in Figure 2.
+The proposed framework employs a multi-level contrastive learning strategy to achieve fine-grained representation learning without the need for class priors, as illustrated in Figure 2.
 Figure 2: Architecture of the proposed Contrastive Disentangling (CD) framework.
@@
   • 4.
-    Feature Predictor: To capture diverse and disentangled feature attributes, the latent representations are further processed by a feature predictor $h(\cdot)$.
+    Feature Predictor: The latent representations are processed by the feature predictor $h(\cdot)$ to capture diverse and disentangled attributes.
 This module, structurally similar to the instance projector but with a sigmoid activation in the final layer, outputs predictions $\mathbf{y}^{(1)}=h(\mathbf{z}^{(1)})$ and $\mathbf{y}^{(2)}=h(\mathbf{z}^{(2)})$.
@@
     Loss Functions: The total loss function of CD integrates three components: the instance-level contrastive loss $\mathcal{L}_{\text{inst}}$, the feature-level contrastive loss $\mathcal{L}_{\text{feat}}$, and the normalized entropy loss $\mathcal{L}_{\text{entropy}}$.
-These losses work synergistically to promote feature diversity, encourage independence among the learned features, and prevent feature collapse.
+These losses work together to ensure that the learned features are diverse, independent, and well-distributed, while avoiding feature collapse.
 3.2 Instance-level Contrastive Learning
-Instance-level contrastive learning aims to learn discriminative features by bringing representations of different augmented views of the same instance closer while pushing apart representations of different instances.
-This dual objective enhances the model’s ability to distinguish between distinct instances and recognize invariant features across different augmentations.
+Instance-level contrastive learning seeks to learn discriminative representations by bringing augmented views of the same instance closer while pushing apart representations of different instances.
+This dual objective improves the model’s ability to differentiate between instances and capture invariant features across augmentations.
 Given an input batch $\mathbf{x}$ of size $N$, two augmented views $\mathbf{x}^{(1)}$ and $\mathbf{x}^{(2)}$ are generated.
@@
 for $i,j\in\{1,2,\dots,2N\}$.
-The purpose of using cosine similarity is to measure how closely related two feature vectors are, with a higher value indicating greater similarity.
+Cosine similarity measures the relationship between latent representations, with higher values indicating greater similarity.
 Building on this similarity measure, we define the instance-level contrastive loss $\mathcal{L}_{\text{inst}}$, which aims to maximize the similarity between positive pairs (i.e., the other augmented view of the same instance) and minimize the similarity between negative pairs (i.e., views of different instances).
@@
 where $\text{pos}(i)$ denotes the index of the positive sample corresponding to $i$, $\tau_{\text{inst}}$ is a temperature parameter controlling the concentration level of the distribution, and $\mathbbm{1}_{[k\neq i]}$ is an indicator function that equals 1 when $k\neq i$ and 0 otherwise.
-This formulation follows the NT-Xent loss used in previous work [2, 17, 18, 19].
 The total instance-level contrastive loss is computed by averaging over all $2N$ samples:
@@
 3.3 Disentanglement Loss
-To ensure that the learned feature representations are both diverse and independent, we introduce two complementary loss functions: the feature-level contrastive loss, which promotes diversity by encouraging each feature head to focus on distinct attributes, and the normalized entropy loss, which ensures that the captured features are meaningful and prevalent across the dataset.
+To ensure that the predicted feature attributes capture meaningful variation across the dataset, we introduce two complementary loss functions: the feature-level contrastive loss, which promotes independence by encouraging each feature head to focus on distinct attributes, and the normalized entropy loss, which ensures that the learned features are well-distributed and representative of the overall data distribution.
 3.3.1 Feature-Level Contrastive Loss
-The feature predictor $h(\cdot)$ produces predictions $\mathbf{y}^{(1)}$ and $\mathbf{y}^{(2)}$, each of size $N\times K$, where $K$ is the number of feature heads.
-To apply feature-level contrastive learning, we concatenate the transposed outputs of the two views:
+The feature predictor $h(\cdot)$ generates predictions $\mathbf{y}^{(1)}$ and $\mathbf{y}^{(2)}$, each of size $N\times K$, where $K$ is the number of feature heads.
+In our framework, each feature head is treated as a distinct “feature instance” that captures different characteristics of the dataset.
+By introducing feature-level contrastive learning, our method uniquely enables each head to capture complementary and independent attributes.
+To maintain consistency in the formulation, we transpose the output matrix of the feature heads and concatenate the outputs from both views:
 $\mathbf{y}=\begin{bmatrix}(\mathbf{y}^{(1)})^{\top}\\ (\mathbf{y}^{(2)})^{\top}\end{bmatrix}=\begin{bmatrix}\mathbf{y}_{1}\\ \mathbf{y}_{2}\\ \vdots\\ \mathbf{y}_{2K}\end{bmatrix}\in\mathbb{R}^{2K\times N},$ (5)
@@
 $\text{sim}_{y}(i,j)=\frac{\mathbf{y}_{i}\cdot\mathbf{y}_{j}}{\|\mathbf{y}_{i}\|\|\mathbf{y}_{j}\|}$ (6)
-The feature-level contrastive loss encourages each feature head to capture distinct attributes by maximizing the similarity between corresponding heads across different views while minimizing the similarity between different heads.
+The feature-level contrastive loss promotes the independence of feature heads by maximizing the similarity between corresponding heads across different views while minimizing the similarity between different heads.
 For feature head $i$, the loss is defined as:
@@
 where $\alpha$ is a hyperparameter that controls the balance between encouraging diversity and ensuring meaningful feature representations.
-By jointly optimizing these two objectives, the feature heads are encouraged to capture both diverse and informative patterns, leading to higher-quality feature disentanglement.
+By simultaneously optimizing these two objectives, the feature heads are encouraged to capture both diverse patterns and informative representations, leading to better feature disentanglement.
 3.4 Total Loss
@@
 $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{inst}}+\mathcal{L}_{\text{disentanglement}}.$ (12)
-By optimizing both the instance and feature-level objectives, our framework achieves a robust balance between learning discriminative instance-level representations and promoting diverse, meaningful feature representations without relying on class priors.
+By jointly optimizing instance and feature-level objectives, our framework achieves a unique balance between learning discriminative representations and ensuring feature diversity, all without relying on class priors.
+This makes our approach especially suited for complex unsupervised learning tasks where class information is scarce or completely unavailable, providing a robust solution for learning fine-grained and diverse representations.
 Algorithm 1 summarizes the proposed method.
@@
 We conducted experiments on four widely-used datasets: CIFAR-10, CIFAR-100 [16], STL-10 [17], and ImageNet-10 [18].
-These datasets serve as benchmarks to rigorously evaluate the effectiveness of our proposed Contrastive Disentangling (CD) framework.
+These datasets serve as benchmarks to rigorously evaluate the effectiveness of our approach.
 Table 1 provides an overview of the dataset splits, sample sizes, and the number of classes.
@@
 Models that utilized class number information during training are marked with *.
-Dataset Metrics K-means AE DeCNN DAE VAE DCGAN CC* CD-128 CD-256
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Dataset CIFAR-10 CIFAR-100-20 STL-10 ImageNet-10
+Metrics NMI ARI ACC NMI ARI ACC NMI ARI ACC NMI ARI ACC
+K-means 0.087 0.049 0.229 0.084 0.028 0.130 0.125 0.061 0.192 0.119 0.057 0.241
+AE 0.239 0.169 0.314 0.100 0.048 0.165 0.250 0.161 0.303 0.210 0.152 0.317
+DeCNN 0.240 0.174 0.282 0.092 0.038 0.133 0.227 0.162 0.299 0.186 0.142 0.313
+DAE 0.251 0.163 0.297 0.111 0.046 0.151 0.224 0.152 0.302 0.206 0.138 0.304
+VAE 0.245 0.167 0.291 0.108 0.040 0.152 0.200 0.146 0.282 0.193 0.168 0.334
+GAN 0.265 0.176 0.315 0.120 0.045 0.151 0.210 0.139 0.298 0.225 0.157 0.346
+DCGAN 0.265 0.176 0.315 0.120 0.045 0.151 0.210 0.139 0.298 0.225 0.157 0.346
+CC* 0.670 0.523 0.684 0.433 0.244 0.431 0.608 0.443 0.594 0.800 0.695 0.846
 […]
@@
 Table 3 compares CD to models that use class number information.
-CD demonstrated strong performance across all metrics, particularly on more complex datasets like STL-10 and ImageNet-10.
-While ARI and ACC were slightly lower on CIFAR-10 and CIFAR-100 compared to CC, this is expected as CD does not use class number information.
-Importantly, CD achieved an ACC of 0.734 with a batch size of 256, surpassing PICA (0.713) despite not using additional unlabeled data.
+CD demonstrated strong performance across all metrics, particularly on more challenging datasets like STL-10 and ImageNet-10.
+While ARI and ACC were slightly lower on CIFAR-10 and CIFAR-100 compared to CC, this is primarily due to the lower complexity of these smaller datasets, which limits the advantage of CD’s ability to disentangle fine-grained features.
+In datasets with more intricate patterns and greater feature diversity, CD’s strengths are more effectively utilized.
+Importantly, CD achieved an ACC of 0.734 with a batch size of 256, surpassing PICA (0.713) despite not using additional unlabeled data, which consists of 100,000 samples.
 Table 3: Final output performance comparison across models.
@@
 Models that utilized class number information during training are marked with *, while those that utilized unlabeled datasets for STL-10 are marked with .
-Dataset Metrics JULE* DEC* DAC* DCCM* PICA* CC* CD-128 CD-256
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Dataset CIFAR-10 CIFAR-100-20 STL-10 ImageNet-10
+Metrics NMI ARI ACC NMI ARI ACC NMI ARI ACC NMI ARI ACC
+JULE* 0.192 0.138 0.272 0.103 0.033 0.137 0.182 0.164 0.277 0.175 0.138 0.300
+DEC* 0.257 0.161 0.301 0.136 0.050 0.185 0.276 0.186 0.359 0.282 0.203 0.381
+DAC* 0.396 0.306 0.522 0.185 0.088 0.238 0.366 0.257 0.470 0.394 0.302 0.527
+DCCM* 0.496 0.408 0.623 0.285 0.173 0.327 0.376 0.262 0.482 0.608 0.555 0.710
+PICA* 0.591 0.512 0.696 0.310 0.171 0.337 0.611 0.531 0.713 0.802 0.761 0.870
+CC* 0.705 0.637 0.790 0.417 0.221 0.421 0.622 0.539 0.670 0.859 0.822 0.893
 […]
@@
 We evaluated the effect of removing the feature prediction head, which reduces the model to an instance-level contrastive learning framework similar to SimCLR.
-As shown in Table 6, the feature prediction head significantly boosts performance, and its removal results in substantial degradation, highlighting the critical role of feature-level contrastive learning in effective disentanglement (see Table 9 for backbone results).
-In comparison to SimCLR under identical training settings, our model demonstrates significant improvements: on the STL-10 dataset, NMI increases by 13.4%, ARI by 23.1%, and ACC by 10.9%; on ImageNet-10, NMI improves by 26.8%, ARI by 110.2%, and ACC by 53.1%.
-These results present that our feature-level contrastive learning approach enables the model to capture more distinct and diverse semantic features, leading to improved clustering and representation quality in the final output.
+As shown in Table 6, the feature prediction head significantly boosts performance, and its removal results in substantial degradation.
+This highlights the critical role of feature-level contrastive learning in effective disentanglement (see Table 9 for backbone results).
+These findings demonstrate that our feature-level contrastive learning approach enables the model to capture more distinct and diverse semantic features, resulting in improved representation quality in the final output.
 Table 6: Effect of feature-level head on model performance.